### v1

# MMiC: Mitigating Modality Incompleteness in Clustered Federated Learning

Abstract
In the era of big data, data mining has become indispensable for uncovering hidden patterns and insights from vast and complex datasets. The integration of multimodal data sources further enhances its potential. Multimodal Federated Learning (MFL) is a distributed approach that enhances the efficiency and quality of multimodal learning, ensuring collaborative work and privacy protection. However, missing modalities pose a significant challenge in MFL, often due to data quality issues or privacy policies across the clients. In this work, we present MMiC, a framework for Mitigating Modality incompleteness in MFL within the Clusters. MMiC replaces partial parameters within client models inside clusters to mitigate the impact of missing modalities. Furthermore, it leverages the Banzhaf Power Index to optimize client selection under these conditions. Finally, MMiC employs an innovative approach to dynamically control global aggregation by utilizing Markovitz Portfolio Optimization. Extensive experiments demonstrate that MMiC consistently outperforms existing federated learning architectures in both global and personalized performance on multimodal datasets with missing modalities, confirming the effectiveness of our proposed solution.
Machine Learning, ICML
1 Introduction
Federated Learning (FL), introduced by McMahan et al. (McMahan et al., 2017), enables multiple participants to collaboratively train a unified model without sharing raw data, thus preserving privacy. This paradigm is particularly well-suited for applications such as medical image analysis, connected vehicles, blockchain and so on(Kaissis et al., 2020; Pokhrel & Choi, 2020; Lu et al., 2020). However, the absence of data sharing in FL introduces significant challenges, including model heterogeneity, communication overhead, and data distribution disparities.
Multimodal Federated Learning (MFL) extends the FL paradigm to multimodal tasks, which leverage the complementary nature of multiple data modalities (e.g., text, images, and audio) to achieve superior performance on large-scale datasets. However, a common challenge in MFL is modal incompleteness, where certain modalities may be missing for some clients due to hardware limitations or data collection constraints. This issue can significantly degrade model performance, especially in heterogeneous scenarios. Existing approaches to tackling missing modalities primarily concentrate on two main aspects: the Data Processing Aspect, which involves generating absent modalities or their representations from the available ones, and the Strategy Design Aspect, which entails devising flexible model architectures capable of adapting to a variable number of available modalities during training, such as attention-based models (Wu et al., 2024). While these methods have shown promise, they are often limited to specific models or datasets. For instance, Yu et al. (Yu et al., 2024) proposed a specialized LSTM-based model to handle incomplete modalities, but such solutions are not easily generalizable to other complex multimodal tasks because they rely heavily on the specific architecture . Moreover, the lack of a unified approach to address modality incompleteness across diverse tasks and datasets further restricts their applicability in broader real-world scenarios. In addition, the current MFL algorithm is limited to information exchange between clients and the server, lacking interaction between clients. For instance, PmcmFL (Bao et al., 2024) compensates for missing modality through server aggregation prototypes. However, due to constraints such as Non-IID data, the aggregated prototypes may not be suitable for all clients.
Clustered Federated Learning (CFL), a specialized form of FL Personalization (Tan et al., 2023), groups clients into clusters based on models, parameters, or datasets. Collaborative training within each cluster reduces the impact of data distribution disparities between clients. Recent advancements in CFL, such as one-shot clustering methods like PACFL (Vahidian et al., 2023), leverage the decomposability of data to identify client clusters efficiently. Each cluster inherently contains clients with similarities in both data and model parameters. We can leverage this characteristic of federated clustering to improve the training scenarios involving missing modalities. This means that the model parameters from clients within the same cluster that have no missing modalities can be used to fill in the parameters of clients with missing modalities. Due to the constraints of communication efficiency within clusters, we can shrink replaced parameters to only one layer most significantly affected by the missing modalities.
Current CFL algorithms predominantly employ random client selection (Ghosh et al., 2019, 2020; Ruan & Joe-Wong, 2022), which ensures fairness over multiple rounds of updates by granting each client an equal opportunity to participate in aggregation. However, we do not expect clients with severe missing modalities to enjoy the same level of aggregation rights. Therefore, by analyzing performance variations across client training rounds, we can determine whether a client is frequently affected by missing modalities. This allows us to adjust the probability of client selection accordingly.
Since the global model, due to the intermediate aggregation process involving the cluster-client hierarchy, cannot directly perceive changes in clients’ missing modalities, we can employ certain functions or metrics to evaluate the training process within each cluster. This evaluation yields a specific value that can be reported to the server, enabling the server to determine the aggregation weights for the current round.
To achieve these goals, we propose MMiC framework that integrates our solutions for the three issues. In particular, we employ the rate of parameter variation to identify the parameters to be replaced. Moreover, the parameters from clients possessing complete modality data and belonging to the same cluster will be utilized to supplant those of clients with incomplete modality data. To enhance the current cluster client selection strategy in the presence of missing modalities, we incorporate the Banzhaf Power Index (BPI), a well-known coalitional game theory, to assess the contribution of each client. This approach enables us to preferentially select the clients who are more contributable, to perform more efficient and effective aggregation processes. We further employ Markovitz Portfolio Optimization, a return-risk balancing asset optimization algorithm to assesses the training process on clusters in a nonlinear manner compared to the standard weighted averaging (Wang & Ji, 2024). The global model adjusts the weights of parameter updates based on the scores obtained from the clusters, thereby reducing the impact of missing modalities in local clients on the global model.
MMiC is model-independent allowing for integration with various multimodal base models. It is also modality-agnostic, enabling its application across diverse multimodal scenarios regardless of the specific modalities involved. Our experiments demonstrate that MMiC significantly outperforms traditional FL and MFL algorithms in terms of both accuracy and stability. In summary, the main contributions of our work are as follows:
  • We propose a parameter substitution mechanism based on clustered federated learning, to mitigate the impact of missing modalities during training.
  • We enhance the random client selection strategy by integrating the Banzhaf Power Index to favor clients with more significant.
  • We are the first to apply the Markowitz Portfolio Optimization theory to evaluate the training on clusters by incorporating the evaluation results into the global aggregation process to mitigate the negative impact from missing modalities to the global server.
2 Related Works
2.1 Clustered Federated Learning (CFL)
Centralized federated learning trains a global model in a central server by aggregating gradients collected across numerous participating clients. The heterogeneity of clients in federated learning introduces challenges such as client drift, feature drift, and label drift, which significantly affect the performance of FL models. Personalized federated learning (PFL) has sprung up designed to better suit the individual characteristics of each client to address the heterogeneity. PFL strategies like Multi-Task Learning, Personalization Layers and Clustered Federated Learning (CFL) have attracted considerable attention in the literature (Smith et al., 2017; Tan et al., 2023; Mansour et al., 2020; Wu et al., 2020).
The fundamental concept of CFL is to group clients based on the similarity of their model parameters, data, or gradients. CFL emphasizes improving cooperation within these clusters to optimize the performance of cluster-specific models, while also enhancing overall global model performance. Ghosh et al. (Ghosh et al., 2019) first introduced the architecture for CFL, and recommended clustering based on the similarity of model parameters. This architecture has been refined through iterative clustering techniques such as the Iterative Federated Clustering Algorithm (IFCA) (Sattler et al., 2021). Fedsoft further advanced the iterative CFL by allowing a client to belong to multiple clusters, outperforming previous efforts. However, these methods incurred substantial resource consumption. One-shot clustering was then introduced by Dennis et al. (Dennis et al., 2021) for more efficient clustering. Their proposed method efficiently determined clustering attribution, identifying the uniqueness of client metrics within complex federated architectures. Vahidian et al. (Vahidian et al., 2023) proposed a one-shot CFL algorithm based on data subspace clustering, They measured the similarity between clients by decomposing the dataset via Singular Value Decomposition (SVD). Our proposed MMiC framework is under the same setting as the one-shot CFL (Vahidian et al., 2023; Yang et al., 2024) and extends it to multimodal scenarios with missing modalities.
2.2 Multimodal Federated Learning (MFL)
Multimodal Federated Learning (MFL) extends standard federated learning to address multimodal datasets on clients and global server.
Challenges and Advances in MFL. FedCMR is the first to address multimodal tasks, though it encountered challenges such as modality discrepancies and missing data. Recent works have made strides in tackling these issues. For instance, Chen and Li (Chen & Li, 2022) proposed MFL-HGB, which adaptively computes optimal modality blending and local model weighting to mitigate overfitting. Similarly, Fed-MP employs adaptive aggregation of local model weights and a multimodal prototyping mechanism to enhance generalization to unseen classes (Zeng et al., 2024). These works collectively provide a foundation for understanding the challenges and solutions in MFL (Che et al., 2023).
MFL with Missing Modality. CACMRN (Xiong et al., 2023) is the first FL framework to propose solutions for missing modalities. They aim to achieve modality reconstruction using existing multi-modal data by training multiple functionally diverse models. Yu et al. (Yu et al., 2023) aim to address the issue of missing modalities in cross-modal retrieval. Clients, each with distinct modalities, learn the representation information of specific modalities and upload it to the server for contrastive learning. Fed-Multimodal (Feng et al., 2023) is a multimodal benchmark built upon the FL algorithm. The work presented a range of models and datasets and proposed a method to initialize the missing modal vector as zero. PmcmFL (Bao et al., 2024) applied the common prototype exchange method of federated learning against Non-IID to the missing modality, and they store prototypes of all modalities for all classes from each client on the server. When encountering missing modalities, it utilizes the corresponding class-specific prototypes to fill in the representations. Additionally, they applied the MIX-UP (Zhang et al., 2018) method in data augmentation into the additional training dataset. In contrast, our approach does not impose restrictions on the types of multimodal tasks, meaning that we do not require the use of specific models.
3 Preliminaries
Before jumping into the technical details of MMiC, we first introduce the fundamental concepts of CFL and the relevant concepts in the game theory and finance domain that have been adopted by MMiC. To assist the description, we list the notations used throughout the paper in Table 1.
Table 1: Main notations.
Notation Description
$K$ Number of clients
$c$ Client
$M$,$m$ Number, index of cluster
$\mathcal{G},\mathcal{M}$ Server, cluster
$T$, $t$ Global Training rounds
$E$, $e$ Local training epochs
$B$, $b$ Batch number, b-th batch
$I,i$, $J,j$ Temporary index or sequence
$\Theta,\theta$ Model or parameters
$\ell$, $L$ Local loss, global Loss
$\eta$ Learning rate
$g$ Gradient
$\mathcal{D}$ Dataset
$R$ Client joining rate each round
$\gamma_{t}$ Portfolio score in $t$-th round
$a,\alpha$ Performance gap
$\sigma$ Risk of federated cluster
$\phi(*)$ Banzhaf interaction index
$w_{k}$ Weight of client $k$
$P$, $\mathcal{P}$ Number of layers, the poverty layer
$\lambda,\epsilon,\tau$ Control constant
$\mathcal{S}$ Subset
3.1 Clustered Federated Learning
Given a global server $\mathcal{G}$ and $K$ clients $\{c_{i}\}^{K}_{1}$, let $\mathcal{D}_{k}$ over $\mathcal{X}*\mathcal{Y}$ be the local dataset of the $k$-th client $c_{k}$, with $\mathcal{X}$ and $\mathcal{Y}$ represent the data inputs and labels, $n_{k}=|\mathcal{D}_{k}|$ be the dataset size. Client weight is calculated as $w_{k}=\frac{n_{k}}{\sum_{k}^{K}n_{k}}$. $c_{k}$ and $\mathcal{G}$ possess the model $\theta_{k}$ and $\theta_{\mathcal{G}}$ respectively and the models are the same when initialized.
Learning Objective. During the FL training process, each cluster model, on the one hand, is regarded as a mesosphere agent that can optimize the training loss from clients within the cluster; and on the other hand, it serves as a local model to the global model, providing stronger generalization ability. In each round of federated training, on the client, it is to learn local model $\theta_{k}$ by minimizing the local empirical loss:
$F_{i}(\theta_{k})=\sum_{\xi\in\mathcal{D}_{k}}\frac{\ell(\theta_{k},\xi)}{n_{k}}$ (1)
where $\ell(\cdot,\cdot)$ is the pre-defined task-driven loss function and $\xi$ represents sampled training data in a batch. The clusters are defined by partitioning $K$ clients into $M$ disjoint subsets. Different from traditional federated selection, clustered federated selection is refined to each cluster, and each round has client subset $\mathcal{S}_{m}$ selected in $\mathcal{M}_{m}$. The weight of cluster $m$ is usually calculated as $w_{m}=\frac{\sum_{k\in\mathcal{M}_{m}}n_{k}}{\sum_{1}^{M}\sum_{k\in\mathcal{M}_{m}}n_{k}}$. Therefore, the global objective function can be expressed as:
$\quad min\quad\sum_{m=1}^{M}\sum_{k\in\mathcal{M}_{m}}w_{k}\ell(\boldsymbol{\theta}_{k})$ (2)
Model Aggregation. A standard way to obtain a cluster model from clients’ models in CFL is to adopt the FedAvg (Li et al., 2020):
$\boldsymbol{\Theta}_{m}=\sum_{k\in\mathcal{M}_{m}}w_{k}\boldsymbol{\theta}_{k}$ (3)
There are various aggregation methods applied to aggregate the global model from clusters’ models. FedAdagrad (Reddi et al., 2021b) exploits dynamic aggregation to combine the clients’ parameter updates from the current round with the stored historical parameter updates:
$\Delta_{m,t}=\boldsymbol{\Theta}_{m,t}-\boldsymbol{\Theta}_{\mathcal{G},t-1}$ (4)
$\boldsymbol{\Delta}_{t}=\frac{1}{M}\sum_{m=1}^{M}w_{m}\boldsymbol{\Delta}_{m,t}$
$\boldsymbol{mtumn}_{t}=\beta\boldsymbol{mtumn}_{t-1}+(1-\beta)\Delta_{t},\beta\in[0,1)$
$\boldsymbol{\upsilon}_{t}=\boldsymbol{\upsilon}_{t-1}+\boldsymbol{\Delta}_{t}^{2}$
$\boldsymbol{\theta}_{\mathcal{G},t}=\boldsymbol{\theta}_{\mathcal{G},t-1}+\eta\frac{\boldsymbol{mtumn}_{t}}{\sqrt{\boldsymbol{\upsilon}_{t}}+\tau}$
where $\Delta_{m,t}$ [^1] Both $\Theta_{m,t}$ and $\Theta_{m}$ represent the cluster model. Here we add the time stamp $t$ to denote $t$-th round of updating and $\boldsymbol{\Delta}_{t}$ is the difference between the current cluster model and the global model. $\boldsymbol{mtumn}_{t}$ is updated iteratively and collects the historical parameter updates, while $\beta$ balances the historical parameter updates $\boldsymbol{mtumn}_{t-1}$ and $\boldsymbol{\Delta}_{t}$. A larger $\beta$ indicates the retention of more historical parameter information. $\boldsymbol{\upsilon}_{t}$ is an accumulated squared change estimate used to measure changes in parameters and adjust the learning rate to improve the stability and convergence speed of the optimization process. MMiC relies on FedAdagrad and derives an extended version (Section 4.3).
3.2 Banzhaf Power Index
The Banzhaf Power Index (BPI) is a game theory metric used to quantify the power of an individual voter in a decision-making body where votes are not necessarily equal (Dubey & Shapley, 1979; Laruelle & Valenciano, 2001). The index is calculated by considering all possible coalitions (i.e., combinations of voters) that can be formed within a voting body and determining the importance of each voter in each coalition. A voter $k$ is considered a core member if their presence in a coalition changes the outcome (denoted as $A_{\mathcal{S}}$) from a loss to a win:
$k\,is\,core?\begin{cases}True&\text{if }A_{\mathcal{S}^{\prime}}\,is\,a\,loss\text{ and }A_{\mathcal{S}}\,is\,a\,win\\ False&\text{otherwise}\end{cases}$ (5)
where $\mathcal{S}^{\prime}$ is the set that excludes voter $k$: $\mathcal{S}^{\prime}=\mathcal{S}\setminus\{k\}$. BPI could be adopted in federated learning settings to help select the most contributable clients (i.e., a subset of the coalition) during the training process, making the client selection more effective.
3.3 Markowitz Portfolio Optimzation
The Markowitz Portfolio Optimization model is an investment optimization model that aims to obtain an optimal investment portfolio with maximized returns and minimized risks (Miller, 1960). The model can assess the current investment process through Risk-Adjusted Return Coefficient RARC ($\gamma$). The quadratic programming model of Markowitz Portfolio Optimization is formulated as follows:
$minimize\quad\gamma=\lambda\bar{\sigma}-(1-\lambda)A$ (6)
indicating RARC is influenced by three variables: investment risk $\bar{\sigma}$, the investment return $A$ and the risk tolerance factor $\lambda$. $\lambda$ signifies the investor’s risk tolerance within the range of [0,1] - When $\lambda$ is set to 0.0, the maximum return value is achieved. $A$ denotes the total portfolio return across all assets. $\bar{\sigma}$ is calculated as follows:
$\bar{\sigma}=\sum_{i=1}^{K}\sum_{j=1}^{K}w_{i}w_{j}\sigma_{ij}$ (7)
$\sigma_{ij}=\frac{1}{T-1}\sum_{t=1}^{T}(\alpha_{i,t}-\bar{\alpha}_{i,t-1})(\alpha_{j,t}-\bar{\alpha}_{j,t-1})$
where $\sigma_{ij}$ is the covariance between assets $i$ and $j$. $w_{i}$ and $w_{j}$ are the weights of assets $i$ and $j$, respectively. $K$ is the number of assets. $\bar{\alpha}_{i,t-1}$ denotes the historical average return of assets $i$ across the preceding $t-1$ rounds. $\sigma_{ij}$, the risk between assets $i$ and $j$ is obtained by subtracting the corresponding average return from each round in the history of the asset. This formulation can be extended to federated learning scenarios, where $\sigma_{ij}$ quantifies the risk or divergence of local models, and $\bar{\alpha_{i,t-1}}$ reflect the historical performance of individual participants in the federated system.
4 Methodology
Figure 1: The MMiC architecture is divided into three layers: Server, Cluster, and Client. The right part illustrates the process of parameter substitution within a cluster. The MPO module primarily operates during the aggregation process from clusters to the global model, while the Banzhaf client selection is applied to the client selection within each cluster.
Figure 1 depicts the architecture of MMiC. The left part illustrates the overall architecture including the clients at the bottom, clusters in the middle, and the global server at the top.
Each cluster accommodates multiple clients, and each client hosts multimodal data with missing modalities locally. The cluster model is derived from the federated average aggregation (Section 3.1). Within each cluster, MMiC handles the missing modality issue and the client selection task. First, a parameter layer substitution mechanism allows clients with missing modalities to obtain knowledge from clients with complete modality data within the same cluster (Section 4.1). Second, client selection is performed leveraging the Banzhaf Power Index, which identifies a set of core clients that contribute the most effectively to the cluster’s model (Section 4.2).
For global aggregation, MMiC extends FedAdagrad by adopting the Markowitz Portfolio Optimization (Section 4.3). This method incorporates client training evaluation results from each cluster and reduces the impact of missing modalities by down-weighing the contribution from clusters that contain missing modalities.
4.1 Cluster-based Parameter Substituion
We are targeting the multimodal scenario when the dataset has both textual $X^{text}$ and image $X^{image}$ data, thus we form the input as a pair {$X^{text}_{i},X^{image}_{i}$} to the $i$-th data in $\mathcal{D}$. The missing modality randomly occurs resulting in data samples in the form of {$X^{null}_{i},X^{image}_{i}$} or {$X^{text}_{i},X^{null}_{i}$} [^2] Please note that MMiC is modality-agnostic – it remains effective regardless of the type or number of modalities, as long as modalities are consistent across clients. We have additional experiments on other modalities shown in Section 5.6.
To mitigate the consequences of missing modalities in each client, parameters that are less impacted by modality incompleteness are used to substitute those that are more significantly affected. We first propose a substitution strategy that can assist these clients in the training rounds. This strategy is inspired by the CFL setting, where a cluster model represents an aggregation of client models within the same cluster. The core idea is to substitute the parameters of client models with missing modality data with those of clients in the same cluster that have complete modality data. This approach involves two key tasks: (i) identifying the parameters to be substituted and (ii) ensuring the efficiency of the method.
Poverty Parameter Identification. In one round of client model training, we identify the layer that is the most affected by missing modalities as $poverty$ layer. We use the parameter difference between the last batch $b-1$ and the current batch $b$ to define the degree of this effect when the missing modality occurs in batch $b$ and batch $b-1$ is with complete modality [^3] We are aware of the missing modality status of the data.:
$\Delta\boldsymbol{\theta}_{k,t}^{e,b}(p)$ $=\text{mean}\left(\left|\frac{\boldsymbol{\theta}_{k,t}^{e,b}(p)-\boldsymbol{\theta}_{k,t}^{e,b-1}(p)}{\boldsymbol{\theta}_{k,t}^{e,b-1}(p)}\right|\right)$
$\text{where}\quad\Delta\boldsymbol{\theta}_{k,t}^{e,b}(p)\in\mathbb{R}$ (8)
where $e$ is the epoch, $t$ is the training round, $k$ is the client. $\Delta\boldsymbol{\theta}_{k,t}^{e,b}(p)$ represents average parameters change rate of the $p$-th layer. As $\Delta\boldsymbol{\theta}_{k,t}^{e,b}$ is a set that contains all the layers change rate. We collect all the batches in the $e$-th epoch that contain complete modality data, while the preceding batch contains missing modality data, and form a batch set $\mathcal{B}*$. Therefore, the relative change rate for the layer $p$ in one epoch could be expressed as:
$\Delta\boldsymbol{\theta}_{k,t}^{e}(p)=\frac{\sum_{b\in\mathcal{B}^{*}}\Delta\boldsymbol{\theta}_{k,t}^{e,b}(p)}{|\mathcal{B}^{*}|}$ (9)
After calculating the average rate of change for all layers in the model in one epoch, we identify the layer with the highest rate of change across all epochs. This layer would be considered as the $poverty$ layer with index $\mathcal{P}_{k,t}$:
$\mathcal{P}_{k,t}=\arg\max_{\begin{subarray}{c}p\in\{1,2,\dots,P\}\end{subarray}}\Delta\boldsymbol{\theta}_{k,t}(p)$ (10)
It is worth mentioning that we select parameters from only one layer to reduce communication consumption.
Poverty Parameter Substitution (PPS). After identifying the poverty parameter in $c_{k}$, we propose to replace these parameters with the parameters from other clients in the cluster that have complete modality data. Specifically, we compute a weighted average of the corresponding parameters of other clients and replace the poverty parameter in $c_{k}$.
$\boldsymbol{\theta}_{k,t}(\mathcal{P}_{k,t})$ $=\sum_{j=1}^{J}w_{j}\boldsymbol{\theta}_{j,t}(\mathcal{P}_{k,t})$
$\quad\text{where }\mathcal{P}_{j,t}=\varnothing,\quad K+J=|\mathcal{M}_{m}|$ (11)
$c_{k}$ represents the poverty client and $c_{j}$ is the client that trained with the complete modality round, and both $c_{k}$ and $c_{j}$ belong to the cluster $\mathcal{M}_{m}$.
Optimizers Could Approximate Parameter Difference. Saving parameters between two batches is resource-intensive and time-consuming. However, we can roughly deduce the rate of change by utilizing the gradient derivation function of the optimizer. Below, we provide the calculation process for parameter variations during the SGD and Adam optimization processes. For SGD, $\eta$ is the learning rate and $\boldsymbol{g}_{k,b}$ is the gradient for $c_{k}$ in $b$ batch.
$\Delta\boldsymbol{\theta}_{k,t}^{e,b}=\frac{(\boldsymbol{\theta}_{k,t}^{e,b-1}-\eta_{k,t}\boldsymbol{g}_{k,b})-\boldsymbol{\theta}_{k,t}^{e,b-1}}{\boldsymbol{\theta}_{k,t}^{e,b-1}}\\ =\frac{-\eta_{k,t}\boldsymbol{g}_{k,b}}{\boldsymbol{\theta}_{k,t}^{e,b}+\boldsymbol{g}_{k,b}}$ (12)
Similarly, we can utilize the Adam optimizer to deduce the average rate of change, here $\hat{\boldsymbol{momentum}}$ is the bias-corrected moment estimates and $\hat{\boldsymbol{\nu}}$ is the second-moment estimate in Adam optimizer (Kingma & Ba, 2015):
$\Delta\boldsymbol{\theta}_{k,t}^{e,b}=\frac{\eta_{k,t}abs(\-\hat{\boldsymbol{momentum}}_{b})}{abs(\boldsymbol{\theta}_{k,t}^{e,b})(\sqrt{\hat{\boldsymbol{\nu}}_{b}}+\epsilon)}$ (13)
4.2 Banzhaf Client Selection in Clusters
The standard procedure for current CFL involves randomly selecting a subset of clients $\mathcal{S}_{m}$ from each cluster for training in each round, followed by a global aggregation. While the random selection is considered fair, there is a possibility that the cluster models aggregated by the selected clients may yield subpar results if some of the selected clients have a significant number of data instances with missing modalities. The negative impact will be transited to the global model through aggregation. Moreover, the negative impact will be enlarged over successive rounds of client training and cluster model aggregation, further degrading the performance of the global model.
Few CFL work addresses this issue and quantifies the impact of client selection for each cluster. We propose a novel method to fill the gap by leveraging Banzhaf Power Index (BPI) to select the core clients. It comprises two main steps. First, we identify core members within each subset selected in each round and tally the times of core members, denoted as core times. Then, by leveraging the historical selection times and core times, we devise a new client selection probability formula, which is applied in our subsequent client selection process. Banzhaf client selection does not require extra communication.
Core Members Identification. At training round $t$, each cluster $\mathcal{M}_{m}$ makes a random client selection. We can obtain the $i$-th client’s local model performance $a_{i,t}$ (i.e., accuracy, F1 score or RSum) of each client in the $t$-th round, and the performance $a_{i,t-1}$ in the ($t$-1)-th round. We calculate the return $\alpha_{i,t}$ of the client for the $t$-th round based on the difference between these two rounds of performances as follows:
$\alpha_{i,t}=a_{i,t}-a_{i,t-1}$ (14)
where $\alpha_{i,t}<0$ indicates that the client is generating a negative return to the investment portfolio; conversely, $\alpha_{i,t}>0$ signifies a positive return. Then, the cluster’s return is calculated as the sum of the client returns:
$A_{\mathcal{S}_{m,t}}=\sum_{i=0}^{|\mathcal{S}_{m,t}|}w_{i}\alpha_{i,t}$ (15)
For each cluster $\mathcal{S}_{m,t}$, we have its total return $A_{\mathcal{S}_{m,t}}$ by Equation (15). Correspondingly, we can calculate the total return for $\mathcal{S}_{m,t}^{{}^{\prime}}$ when excluding client $c_{i}$, $\𝑖\mathcal{S}_{m,t}^{\prime}=\mathcal{S}_{m,t}\textbackslash\{i\}$: Then, we use the return of cluster model to evaluate the core members:
$c_{i}\,is\,core\,member?\begin{cases}True&A_{\mathcal{S}_{m,t}}\geq\alpha^{m}_{t}\text{ and }A_{\mathcal{S}_{m,t}^{{}^{\prime}}}<\alpha^{m}_{t}\\ False&\text{otherwise}\end{cases}$ (16)
The above equation describes the process of determining whether client $c_{i}$ is a core member during the $t$-th round of client random selection. We utilize a counter $\phi(*)$ that starts with 0 to tally the number of times each client has been identified as a core member:
$\phi(i)=\sum_{t=1}^{T}\mathbb{I}\left(c_{i}\,is\,a\,core\,member\,\right)$ (17)
where $\mathbb{I}(*)$ is the indicator function that returns 1 if the condition * is satisfied, otherwise returns 0.
Client Selection Probability. We could design a method to measure the probability of a client being selected in the current round based on its core times. Our goal is to ensure that a higher core time corresponds to a higher probability. However, this way will make clients with small core times unlikely to be selected later. To solve this problem, we set another counter $\mathcal{T}(*)$ that records the number of times the client has been selected in its history, and we determine the final selection probability by balancing the historically selected times $\mathcal{T}(i)$ and the core times $\phi(i)$:
$prob(i)=\frac{e^{\frac{\tau\phi(i)}{\mathcal{T}(i)}}}{\sum_{i\in\mathcal{M}_{m}}e^{\frac{\tau\phi(i)}{\mathcal{T}(i)}}}$ (18)
For client $i$ in cluster $\mathcal{M}_{m}$, its probability in the $t$-th round to be selected is $prob(i)$. We increase the disparity in probability allocation for each client through an exponential function, where $\tau$ is a temperature constant. As $\tau$ increases, the disparity in probability increases, and as $\tau$ decreases, the disparity in probability decreases.
4.3 Markowitz Profolio Optimization for Global Aggregation
As introduced in Section 3.3, Markowitz Portfolio Optimization is a framework for constructing an investment portfolio that aims to maximize returns for a given level of risk. We are the first to integrate this theory into CFL by considering to mitigate the negative effects of local client models (e.g., those with missing modalities) on the global model. Specifically, we propose a mechanism where, in each communication round, the performance of the selected clients within a cluster $\mathcal{M}_{m}$ is quantified through a non-linear scoring function. This score serves as a direct indicator of the cluster’s performance during the round, allowing for adaptive control over the magnitude of the global model’s updates. By leveraging this approach, we dynamically regulate the contribution of each cluster to the global model aggregation, thereby reducing the impact of underperforming or unreliable clients while maintaining the global model’s robustness.
To achieve this, we require three core indicators to construct our “investment portfolio” for the cluster $\mathcal{M}m$: the return (i.e., payoff) of each cluster ($A_{m,t}$, Equation (15)); the risk associated with the cluster ($\bar{\sigma}_{m,t}$); and the tolerance for return-risk ratio for each cluster ($\lambda$).
Cluster Risk. According to the asset risk calculation formula Equation (7), we obtain the covariance as follows:
$\bar{\sigma}_{m,t}$ $=\sum_{i=1}^{|\mathcal{S}_{m}|}\sum_{j=1}^{|\mathcal{S}_{m}|}w_{i}w_{j}\Bigg{\{}\frac{1}{T-1}\sum_{t=1}^{T}$
$\quad(\alpha_{i,t}-\bar{\alpha}_{i,t-1})(\alpha_{j,t}-\bar{\alpha}_{j,t-1})\Bigg{\}}$ (19)
where $\bar{\alpha}_{i,t-1}$ represents the historical average return for client $i$ in round $t-1$. $w_{i}$ and $w_{j}$ are weights assigned to the $i$-th and $j$-th clients respectively. Please refer to Section 3.1 for weights’ calculation. Then the cluster risk is formulated as follows according to Equation (7):
$\quad\gamma_{m,t}=\lambda\bar{\sigma}_{m,t}-(1-\lambda)A_{m,t}$ (20)
where $\lambda$ [0,1] is the risk tolerance constant. We defaultly set it as 0.5 to balance risk and return. Our goal is to minimize the sum of weighted cluster RARC (cluster weight calculation is introduced in Section 3.1):
$minimize\quad\gamma_{t}=\sum_{m=1}^{M}w_{m}\gamma_{m,t}$ (21)
We analyzed how we achieved the time complexity of MPO in Appendix D.
Participate in Global Aggregation. A higher $\gamma_{t}$ signifies that the clusters have undergone a relatively poor local client training process. In such scenarios, our objective is to diminish the impact of the current round’s parameter updates on the global aggregation by down-weighting the cluster parameters associated with higher $\gamma$. To accomplish this, we introduce a dynamic $\beta^{*}$ to replace the static $\beta$ in FedAdagrad, resulting in FedAdagradMPO, which effectively captures the influence of $\gamma_{t}$:
$\beta^{*}=\beta+\frac{tanh(ReLU(\gamma_{t}))}{\tau}$ (22)
When $\gamma_{t}>0$, then $\beta^{*}>\beta$, which helps retain more historical parameter updates in the global aggregation. Otherwise, $\beta^{*}=\beta$. Here $\tau$ serves as the smoothness control constant for the range. We suggest setting the temperature to range [10,20], and it depends on the value set for $\beta$.
4.4 Variants of MMiC
We propose two MMiC variants by extending two existing methods that address modality incompleteness. Specifically, by integrating these methods with the three proposed modules in Section 4.1 - 4.3, we form the two MMiC variants.
MMiC (Modality Reconstruction). In this variant, we adopt the reconstruction idea from  (Saharia et al., 2022) and incorporate two modality reconstruction models, Text-to-Image and Image-to-Text, at each cluster to reconstruct the missing modality using the data from the complete modality.
MMiC (GMC-Fusion). The other MMiC variant is based on the GMC-Fusion model (Poklukar et al., 2022). This variant necessitates the additional projecting of information from two modalities into the shared space, followed by computing the Geometric Multimodal Contrastive loss using all available modality representations and shared space representations from a batch. In this way, the missing modality information could be filled in the shared space.
5 Experiments
Table 2: Global and personalized performance compared with different algorithms. Bold with * indicated the best performance, while bold black indicated excellent or stable top three performance. Since the PmcmFL algorithm is only for classification tasks, it can’t be compared in Flick30k. MR stands for Modality Reconstruction and GF is GMC-Fusion.
Global CrisisMMD (F1 score) UPMC-Food101 (Accuracy) Flickr30k (RSum score)
mm@0.2 mm@0.2 mm@0.5 mm@0.2 mm@0.2 mm@0.5 mm@0.2 mm@0.2 mm@0.5
mc@0.3 mc@0.5 mc@0.5 mc@0.3 mc@0.5 mc@0.5 mc@0.3 mc@0.5 mc@0.5
FedAvg [2017] 21.886 ± 2.87 20.531 ± 1.9 24.095 ± 0.178 88.053 ± 1.03 87.626 ± 1.34 88.283 ± 0.83 14.324 ± 1.68 4.704 ± 0.38 3.812 ± 0.16
FedOpt [2021] 29.320 ± 1.5 29.750 ± 2.04 26.292 ± 1.91 88.270 ± 0.53 89.064 ± 1.09 88.324 ± 0.85 16.138 ± 1.06 5.044 ± 0.61 12.576 ± 1.28
FedSoft [2022] 11.371 ± 0.02 8.180 ± 0.12 10.464 ± 0.01 87.471 ± 0.38 85.984 ± 0.79 81.461 ± 0.14 1.189 ± 0.01 1.297 ± 0.01 1.106 ± 0.02
PACFL [2023] 23.239 ± 3.35 20.274 ± 2.18 23.131 ± 3.029 88.152 ± 1.69 88.343 ± 1.31 88.439 ± 0.91 10.168 ± 1.14 8.013 ± 0.79 3.678 ± 0.25
CreamFL [2023] 22.766 ± 0.09 18.131 ± 1.0 16.958 ± 2.93 87.539 ± 0.26 87.102 ± 0.17 87.390 ± 0.88 13.965 ± 0.04 13.857 ± 0.06 11.839 ± 0.02
PmcmFL [2024] 24.610 ± 1.49 22.471 ± 0.44 22.946 ± 0.89 87.916 ± 0.90 85.159 ± 0.13 86.873 ± 0.52 - - -
MMiC *29.462 ± 0.36 27.815 ± 1.44 *28.480 ± 1.86 88.777 ± 0.86 *89.436 ± 1.21 *89.428 ± 0.78 14.266 ± 1.20 8.322 ± 0.177 14.220 ± 1.28
MMiC (MR) 29.108 ± 0.86 *30.471 ± 0.77 28.354 ± 1.36 *88.997 ± 0.33 88.519 ± 2.31 89.044 ± 0.76 10.907 ± 1.21 10.152 ± 0.87 8.614 ± 0.84
MMiC (GF) 27.366 ± 2.22 26.229 ± 4.15 18.354 ± 2.06 88.465 ± 0.539 87.973 ± 1.26 89.056 ± 1.17 *18.333 ± 1.172 *25.418 ± 1.90 *22.707 ± 1.42
Personalized
FedAvg [2017] 11.133 ± 0.49 9.671 ± 0.368 13.454 ± 0.87 77.039 ± 2.69 76.080 ± 1.47 75.786 ± 1.76 34.625 ± 2.39 17.625 ± 0.76 17.907 ± 0.24
FedOpt [2021] 17.417 ± 1.24 18.825 ± 0.16 15.205 ± 1.07 76.073 ± 2.05 76.492 ± 1.40 76.304 ± 2.76 40.275 ± 2.22 33.569 ± 1.89 43.152 ± 1.43
FedSoft [2022] 10.814 ± 0.384 8.012 ± 0.57 8.180 ± 0.01 75.547 ± 0.94 75.108 ± 1.21 73.918 ± 1.41 7.812 ± 1.32 6.103 ± 0.36 3.695 ± 2.52
PACFL [2023] 13.837 ± 1.72 10.454 ± 0.56 12.985 ± 1.97 83.805 ± 1.76 82.915 ± 1.75 83.722 ± 1.96 18.348 ± 0.62 20.171 ± 3.49 16.266 ± 1.08
CreamFL [2023] 11.169 ± 2.37 11.853 ± 0.815 11.338 ± 0.647 78.610 ± 0.411 78.484 ± 4.85 77.008 ± 2.07 76.900 ± 3.88 79.461 ± 1.92 75.122 ± 3.01
PmcmFL [2024] 17.010 ± 0.32 14.662 ± 1.36 14.048 ± 0.06 83.031 ± 1.20 81.673 ± 1.59 79.987 ± 0.82 - - -
MMiC *19.936 ± 0.91 18.911 ± 1.13 18.259 ± 1.62 83.576 ± 0.75 *84.022 ± 1.19 84.057 ± 1.73 90.726 ± 3.28 75.814 ± 3.35 78.979 ± 2.38
MMiC (MR) 19.492 ± 0.27 19.174 ± 1.32 18.101 ± 1.08 83.519 ± 2.31 83.055 ± 3.74 82.948 ± 3.59 41.824 ± 6.70 33.388 ± 4.52 31.489 ± 4.17
MMiC (GF) 19.625 ± 4.797 *22.796 ± 6.04 *18.450 ± 2.73 *84.715 ± 3.13 83.639 ± 2.10 *84.601 ± 4.69 *112.506 ± 5.10 *160.99 ± 6.082 *124.414 ± 2.86
5.1 Datasets, Models and Metrics
Our evaluation of MMiC is based on two multimodal classification datasets and one multimodal retrieval dataset. Because each dataset has distinct characteristics, we followed (Feng et al., 2023; Jeni et al., 2013) in employing different base models and evaluation metrics for each dataset. We also provide learning configurations in Appendix A. CrisisMMD (Ofli et al., 2020). It is a multi-modal dataset for classification consisting of 18,126 samples annotated tweets and images. We brought in a lightweight multi-modal model that used pre-trained embedding layers for input and concatenates shared representations (Gupta et al., 2024; Feng et al., 2023). Following (Feng et al., 2023), we employed the F1 (Goutte & Gaussier, 2005) score to evaluate the performance of CrisisMMD.
UPMC-Food101 (Gallo et al., 2020). This dataset was sourced from Food101, containing 101,000 image-text pairs across 101 food categories. We used CLIP with the official ViT version of the backbone pre-trained model (Radford et al., 2021). We froze the first 11 layers of attention while retaining the last layer of attention and the final output linear layer for parameter updates. We employed accuracy to evaluate its performance.
Flickr30k (Young et al., 2014). It is an image captioning dataset for retrieval, which is comprised of 31,783 images and the corresponding captions. We used SCAN (Stacked Cross Attention Network) (Lee et al., 2018) as the multimodal retrieval model. SCAN is a model designed for image-text matching. We relied on the RSum as the evaluation metric to measure the performance of the retrieval task (Faghri et al., 2018).
5.2 Benchmark Algorithms
We compared MMiC with selected representative works and state-of-the-art works.
FedAvg (McMahan et al., 2017) represented a basic federated architecture where multiple clients collaborated with a central server. FedOpt (Reddi et al., 2021a) proposed three specialized federated aggregation methods, building upon FedAvg. FedSoft (Ruan & Joe-Wong, 2022), as a soft clustering method, introduced a technique that permitted a client to belong to multiple clusters. PACFL (Vahidian et al., 2023), an advanced CFL methodology, employed Singular Value Decomposition (SVD) to deconstruct the dataset into its constituent subspaces, subsequently applying hierarchical clustering techniques to these subspaces for effective clustering. CreamFL (Yu et al., 2023), was a MFL algorithm that employed inter-modal contrastive learning to address the issue of modality imbalance. PmcmFL (Bao et al., 2024), was another MFL for addressing missing modality. They utilized prototype aggregation to address the issue of missing modalities in classification tasks.
5.3 Settings
Cluster and Clients Settings. In this experiment, we employed two distance metrics for clustering: Locality Sensitive Hashing (LSH) sketch (Yang et al., 2024; Liu et al., 2023) and SVD decomposition (Vahidian et al., 2023). All experiments used the LSH sketch for clustering unless otherwise noted. Each cluster had a test set visible to all clients within the cluster. Additionally, the global server provided accessibility to the test set for all clients. We simulated the Non-IID scenario by assigning a label probability table to each client utilizing the Dirichlet distribution (Teh, 2010) with $\beta_{diri}=0.5$. For all the experiments, we set the number of clients to 50 so that ensured each client had at least 200 training data. Our client selection strategy comprised an initial phase of 100 rounds with random selection, followed by a subsequent phase of 200 rounds utilizing Banzhaf client selection (300 global rounds).
Missing Modality Settings. Our setting with missing modalities was governed by two variables: clients with missing modality (mc) and missing modality rate (mm), denoting the proportion of all clients selected to have missing modalities and the proportion of missing samples on these clients, respectively. Our main experiment (Section 5.4) offered three levels of modality incompleteness settings: (mm@0.2, mc@0.3), (mm@0.2, mc@0.5) and (mm@0.5, mc@0.5). each client has a fixed number of missing samples, determined by the mm rate. In each round, among the clients selected for aggregation, we use mc to decide whether the selected clients will activate missing data. Since the selected clients use a random training subset for training in each round, the maximum missing modality rate for a client in that round is $min(\frac{missing\ samples}{subset\ samples},1)$.
5.4 Overall Performance
We evaluated MMiC based on global performance (gbp) and personalized performance (plp) and presented the results in Table 2. To do the evaluation, we made use of the global test dataset to evaluate the global model performance following  (McMahan et al., 2017), and adopted the weighted average performance from client models to evaluate the plp following  (Fallah et al., 2020).
MMiC and its two variants secure the top positions across all tasks and all missing modality scenarios. While MMiC (MR) excelled in classification tasks, being the only algorithm to achieve an F1 score exceeding 30 on CrisisMMD (mm@0.2, mc@0.5) at gbp. MMiC (GF) significantly outperformed all other algorithms in modality retrieval tasks across all three metrics. These results highlighted the superiority of MMiC in handling diverse missing modality scenarios.
FedOpt demonstrated strong gbp on the CrisisMMD dataset (mm@0.2, mc@0.3) and (mm@0.2, mc@0.5). FedSoft’s performance is medium across experiments with multiple missing modalities settings and datasets both in global and personalized evaluations. As a CFL algorithm, PACFL do not exhibit outstanding performance in gbp. In contrast, it demonstrated excellent performance in personalized evaluation. CreamFL is primarily an algorithm designed for retrieval tasks. The overall performance of CreamFL is relatively mediocre except for Flickr30k (mm@0.2, mc@0.5), their method ranked second in this setting both in global and personalized evaluations. PmcmFL currently supports only classification tasks; however, its performance in classification tasks is somewhat limited.
(a) only PPS
(b) w/o PPS
(c) only BPI
(d) w/o BPI
(e) only MPO
(f) w/o MPO
Figure 2: Ablation study under CrisisMMD (mm@0.5 mc@0.5).
5.5 Ablation Study
We conducted ablation experiments of MMiC on CrisisMMD. The results are shown in Figure 2. A further ablation study on Food101 and Flickr30k is provided in Appendix B. BPI (b), MPO (p) and PPS (s) represent Banzhaf Power Selection, Markowitz Portfolio Optimization, and Poverty Parameter Substitution, respectively. The figures show that MMiC consistently achieves higher accuracy in the final rounds across multiple ablation tests. Specifically, both ‘w/o PPS’ (Figure 2(b)) and ‘w/o BPI’ (Figure 2(d)) saw a decrease of more than 5% compared with MMiC. The final performance of ‘w/o MPO’ (Figure 2(f)) is close to MMiC, yet it remained approximately 2% lower than MMiC. In addition, removing any of the two contributions leads to varying degrees of performance decline, with the resulting performance consistently falling between that of MMiC and the configuration with a single module removed (Figure 2(a), 2(c), 2(e)). We conclude that removing any module leads to performance degradation. We further proved the effectiveness of BPI in Appendix C.
5.6 Experiments on other modalities
We further experimented on two additional multimodal datasets with modalities other than image and text, aiming to show that MMiC is modality-agnostic. UCI-HAR (Reyes-Ortiz et al., 2013) and KU-HAR (Sikder & Nahid, 2021) both encompass smartphone sensor data utilized for the classification of diverse human activities. They include accelerometer and gyroscope readings gathered from a multitude of participants. We compared MMiC with FedOpt, which exhibited the good performance in Section 5.4 , for this analysis. The experimental setup involved 30 clients running 150 global rounds, with all other parameters remaining consistent with those outlined in Section 5.3. The findings of this experiment are presented in Table 3. MMiC demonstrated a pronounced superiority over FedOpt, with enhancements of 2% and 6% in the KU-HAR and UCI-HAR respectively.
Table 3: Performance comparison of MMiC and FedOpt on UCI-HAR and KU-HAR datasets.
Dataset/Algorithm MMiC (%) FedOpt (%)
UCI-HAR 86.99 80.89
KU-HAR 75.83 73.24
6 Conclusions
In this work, we introduce MMiC, a model-independent and modality-agnostic framework designed to mitigate the impact of missing modalities in multimodal federated learning. By substituting parameters within clustered federated learning, adapting the Banzhaf Power Index for client selection, and extending Markowitz Portfolio Optimization for global aggregation, MMiC demonstrates improved performance over existing federated architectures in addressing missing modalities. Moving forward, we plan to further enhance the framework’s robustness by substituting additional poverty parameters and exploring more efficient solutions.
Appendix A Learning Settings
In all of our experiments, we set a global learning rate (0.1, 0.01, 0.05) in the three datasets (CrisisMMD, Food101, Flickr30k) respectively to global optimizers ($FedAdagrad_{MPO}$, $FedAdagrad$). Our Parameter difference calculation process is derived from Equation 13 and we used the Adam optimizer with a learning rate of 0.005 for the clients. These parameters were obtained through our simple tuning in the early stage of our experiment. The batch size for the CrisisMMD dataset was set to 15, and 7 for the UPMC-Food101 dataset, 20 for the Flickr30k dataset. All the local epoch was set to 3. The loss function on the client was cross-entropy, except for MMiC (GMC-Fusion), where the loss consists of two parts: cross-entropy and contrastive loss between existing modalities (Poklukar et al., 2022).
Appendix B Ablation Study on Food101 and Flickr30k
We conducted an ablation study on the Food101 and Flickr30k datasets under the settings of mm@0.2 and mc@0.5. The modules are denoted as ‘b’, ‘p’, and ‘s’ (BPI, MPO and PPS respectively following Section 5.5). The full MMiC framework demonstrated robust performance, achieving an accuracy exceeding 90% on Food101 and a score of 11 on Flickr30k under the specified conditions. In contrast, the exclusion of all modules lead to a significant decline in accuracy, reducing it to 86% on Food101 and 4.3 on Flickr30k. The removal of a single module generally yields superior performance compared to the exclusion of two modules, as presented in Table 4.
Table 4: Ablation study on Food101 and Flickr30k under (mm@0.2, mc@0.5).
Configuration Food101 (Accuracy) Flickr30k (RSum)
w/o (b, p, s) 86.97 4.28
w/o s 88.54 6.99
w/o p 88.91 8.18
w/o b 87.26 8.10
only s 88.03 7.08
only p 87.49 4.84
only b 87.41 5.36
MMiC 90.12 11.02
Figure 3: The number of additional training rounds needed for the Random Client Selection to achieve the same performance as the Banzhaf Client Selection (Flickr30k).
Appendix C Effectiveness of Banzhaf Client Selection
The federated architecture had shown the feasibility of client selection, as demonstrated in (Nishio & Yonetani, 2019; Cho et al., 2020). The Banzhaf client selection strategy aimed to optimize global aggregation efficiency by minimizing communication rounds. Figure 3 illustrates the comparison of communication rounds and performance between the setting with Banzhaf client selection and random client selection across three levels of missing modality. We used Flickr30k dataset in this evaluation. It should be noted that randomly selected patterns required additional (12, 24, or 16) training rounds to surpass or approach the performance of the Banzhaf client selection.
Appendix D Time Complexity of MPO
Since the Banzhaf Client Selection do not incur an additional cost, the time complexity of MMiC compared to other works depended on the complexity of Markowitz Portfolio Optimization. The aggregation process of $FedAdagrad_{MPO}$ involves the computation of RARC in each cluster, which is then uploaded to the server. The server utilizes the RARC of each cluster to dynamically adjust the beta parameter for aggregation. Since the computational complexity of beta is linear and negligible, the overall complexity of MPO is primarily dominated by the RARC computation. The time complexity of MPO is $O(\frac{K^{2}}{M}\cdot T+K\cdot T)$. Here we deduce:
  • We got average $\frac{K}{M}$ clients for each cluster. Time complexity of $\sigma_{ij}$ was $O(T)$.
    Time complexity of risk $\bar{\sigma}$:
    $\bar{\sigma}=\sum_{i=1}^{\frac{K}{M}}\sum_{j=1}^{\frac{K}{M}}w_{i}w_{j}\sigma_{ij}.$
    could be obtained from:
    $O\left((\frac{K}{M})^{2}\right)\times O(T)=O\left((\frac{K}{M})^{2}\cdot T\right)$
    In the covariance $\sigma_{ij}$ section, we can employ the concept of prefix sums (Blelloch, 1990) to efficiently store historical data. The time complexity of constructing prefix sums was extra $O(\frac{K^{2}}{M^{2}})$. Hence this part could be improved to $O(\frac{K^{2}}{M^{2}})+O(\frac{K^{2}}{M^{2}})$.
  • Return $A$ computation is to do the weighted sum for each cluster, and the time complexity of $A$ is $O(\frac{K}{M})$ to each cluster.
  • $M$ clusters will upload their RARC to the server and calculate the weighted sum:
    $\left(O(\frac{K^{2}}{M^{2}}+\frac{K}{M})+O(\frac{K^{2}}{M^{2}})\right)\times M=O(\frac{K^{2}}{M}+K)+O(\frac{K^{2}}{M})$
    Meanwhile, MMiC has a total $T$ global rounds. As prefix sums only build at once, we got:
    $O(\frac{K^{2}}{M}+K)\times T+O(\frac{K^{2}}{M})=O\left(\frac{K^{2}\cdot T}{M}+K\cdot T+\frac{K^{2}}{M}\right)$
    $O(\frac{K^{2}}{M})$ was less than $O(\frac{K^{2}\cdot T}{M})$ apparently. Therefore, the time complexity of MPO is $O(\frac{K^{2}}{M}\cdot T+K\cdot T)$.

### v3

@@
 MMiC: Mitigating Modality Incompleteness in Clustered Federated Learning
-Abstract
+Abstract.
 In the era of big data, data mining has become indispensable for uncovering hidden patterns and insights from vast and complex datasets.
@@
 Furthermore, it leverages the Banzhaf Power Index to optimize client selection under these conditions.
-Finally, MMiC employs an innovative approach to dynamically control global aggregation by utilizing Markovitz Portfolio Optimization.
+Finally, MMiC employs an innovative approach to dynamically control global aggregation by utilizing Markowitz Portfolio Optimization.
 Extensive experiments demonstrate that MMiC consistently outperforms existing federated learning architectures in both global and personalized performance on multimodal datasets with missing modalities, confirming the effectiveness of our proposed solution.
-Machine Learning, ICML
-1 Introduction
+Our code released at https://github.com/gotobcn8/MMiC.
+Federated Learning, Multimodal Learning, Missing Modality
+journalyear: 2025copyright: acmlicensedconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Korea.booktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreaisbn: 979-8-4007-2040-6/2025/11doi: 10.1145/3746252.3761140ccs: Computing methodologies Distributed computing methodologiesccs: Information systems Data miningccs: Computing methodologies Distributed artificial intelligenceccs: Computing methodologies Machine learning
+1. Introduction
 Federated Learning (FL), introduced by McMahan et al.
 (McMahan et al., 2017), enables multiple participants to collaboratively train a unified model without sharing raw data, thus preserving privacy.
-This paradigm is particularly well-suited for applications such as medical image analysis, connected vehicles, blockchain and so on(Kaissis et al., 2020; Pokhrel & Choi, 2020; Lu et al., 2020).
+This paradigm is particularly well-suited for applications such as medical image analysis, connected vehicles, blockchain and so on(Kaissis et al., 2020; Pokhrel and Choi, 2020; Lu et al., 2020).
 However, the absence of data sharing in FL introduces significant challenges, including model heterogeneity, communication overhead, and data distribution disparities.
@@
 We can leverage this characteristic of federated clustering to improve the training scenarios involving missing modalities.
-This means that the model parameters from clients within the same cluster that have no missing modalities can be used to fill in the parameters of clients with missing modalities.
+This means that the model parameters from clients within the same cluster that have full modalities can be used to fill in the parameters of clients with missing modalities.
 Due to the constraints of communication efficiency within clusters, we can shrink replaced parameters to only one layer most significantly affected by the missing modalities.
-Current CFL algorithms predominantly employ random client selection (Ghosh et al., 2019, 2020; Ruan & Joe-Wong, 2022), which ensures fairness over multiple rounds of updates by granting each client an equal opportunity to participate in aggregation.
+Current CFL algorithms predominantly employ random client selection (Ghosh et al., 2019, 2020; Ruan and Joe-Wong, 2022), which ensures fairness over multiple rounds of updates by granting each client an equal opportunity to participate in aggregation.
 However, we do not expect clients with severe missing modalities to enjoy the same level of aggregation rights.
@@
 This evaluation yields a specific value that can be reported to the server, enabling the server to determine the aggregation weights for the current round.
-To achieve these goals, we propose MMiC framework that integrates our solutions for the three issues.
+To achieve these goals, we propose MMiC, a framework that integrates our solutions for the three issues.
 In particular, we employ the rate of parameter variation to identify the parameters to be replaced.
@@
 This approach enables us to preferentially select the clients who are more contributable, to perform more efficient and effective aggregation processes.
-We further employ Markovitz Portfolio Optimization, a return-risk balancing asset optimization algorithm to assesses the training process on clusters in a nonlinear manner compared to the standard weighted averaging (Wang & Ji, 2024).
+We further make use of Markowitz Portfolio Optimization, a return-risk balancing asset optimization algorithm to assess the training process on clusters in a nonlinear manner compared to the standard weighted averaging (Wang and Ji, 2024).
 The global model adjusts the weights of parameter updates based on the scores obtained from the clusters, thereby reducing the impact of missing modalities in local clients on the global model.
@@
   • We enhance the random client selection strategy by integrating the Banzhaf Power Index to favor clients with more significant.
-  • We are the first to apply the Markowitz Portfolio Optimization theory to evaluate the training on clusters by incorporating the evaluation results into the global aggregation process to mitigate the negative impact from missing modalities to the global server.
-2 Related Works
-2.1 Clustered Federated Learning (CFL)
+  • We are the first to apply the Markowitz Portfolio Optimization theory to evaluate the training process by incorporating the evaluation results into the global aggregation process to mitigate the negative impact of missing modalities on the global server.
+2. Related Works
+2.1. Clustered Federated Learning (CFL)
 Centralized federated learning trains a global model in a central server by aggregating gradients collected across numerous participating clients.
-The heterogeneity of clients in federated learning introduces challenges such as client drift, feature drift, and label drift, which significantly affect the performance of FL models.
 Personalized federated learning (PFL) has sprung up designed to better suit the individual characteristics of each client to address the heterogeneity.
@@
 (Ghosh et al., 2019) first introduced the architecture for CFL, and recommended clustering based on the similarity of model parameters.
-This architecture has been refined through iterative clustering techniques such as the Iterative Federated Clustering Algorithm (IFCA) (Sattler et al., 2021).
 Fedsoft further advanced the iterative CFL by allowing a client to belong to multiple clusters, outperforming previous efforts.
@@
 Our proposed MMiC framework is under the same setting as the one-shot CFL (Vahidian et al., 2023; Yang et al., 2024) and extends it to multimodal scenarios with missing modalities.
-2.2 Multimodal Federated Learning (MFL)
+2.2. Multimodal Federated Learning (MFL)
 Multimodal Federated Learning (MFL) extends standard federated learning to address multimodal datasets on clients and global server.
@@
 FedCMR is the first to address multimodal tasks, though it encountered challenges such as modality discrepancies and missing data.
-Recent works have made strides in tackling these issues.
-For instance, Chen and Li (Chen & Li, 2022) proposed MFL-HGB, which adaptively computes optimal modality blending and local model weighting to mitigate overfitting.
-Similarly, Fed-MP employs adaptive aggregation of local model weights and a multimodal prototyping mechanism to enhance generalization to unseen classes (Zeng et al., 2024).
+Recent works have made strides in tackling these issues (Zong et al., 2021).
+Graph-based FL represents a special case of CFL, they could be used to handle the heterogeneous modality fusion.
+FedMSplit (Chen and Zhang, 2022) employs a dynamic and multi-view graph structure to adaptively capture the correlations amongst multimodal local models.
+AGHG (Qi and Li, 2024) creates a global knowledge enhancer to extract information from different modalities.
 These works collectively provide a foundation for understanding the challenges and solutions in MFL (Che et al., 2023).
@@
 When encountering missing modalities, it utilizes the corresponding class-specific prototypes to fill in the representations.
-Additionally, they applied the MIX-UP (Zhang et al., 2018) method in data augmentation into the additional training dataset.
 In contrast, our approach does not impose restrictions on the types of multimodal tasks, meaning that we do not require the use of specific models.
-3 Preliminaries
-Before jumping into the technical details of MMiC, we first introduce the fundamental concepts of CFL and the relevant concepts in the game theory and finance domain that have been adopted by MMiC.
-To assist the description, we list the notations used throughout the paper in Table 1.
-Table 1: Main notations.
-Notation Description
-$K$ Number of clients
-$c$ Client
-$M$,$m$ Number, index of cluster
-$\mathcal{G},\mathcal{M}$ Server, cluster
-$T$, $t$ Global Training rounds
-$E$, $e$ Local training epochs
-$B$, $b$ Batch number, b-th batch
-$I,i$, $J,j$ Temporary index or sequence
-$\Theta,\theta$ Model or parameters
-$\ell$, $L$ Local loss, global Loss
-$\eta$ Learning rate
-$g$ Gradient
-$\mathcal{D}$ Dataset
-$R$ Client joining rate each round
-$\gamma_{t}$ Portfolio score in $t$-th round
-$a,\alpha$ Performance gap
-$\sigma$ Risk of federated cluster
-$\phi(*)$ Banzhaf interaction index
-$w_{k}$ Weight of client $k$
-$P$, $\mathcal{P}$ Number of layers, the poverty layer
-$\lambda,\epsilon,\tau$ Control constant
-$\mathcal{S}$ Subset
-3.1 Clustered Federated Learning
+3. Clustered Federated Learning
+Optimization Objective.
 Given a global server $\mathcal{G}$ and $K$ clients $\{c_{i}\}^{K}_{1}$, let $\mathcal{D}_{k}$ over $\mathcal{X}*\mathcal{Y}$ be the local dataset of the $k$-th client $c_{k}$, with $\mathcal{X}$ and $\mathcal{Y}$ represent the data inputs and labels, $n_{k}=|\mathcal{D}_{k}|$ be the dataset size.
-Client weight is calculated as $w_{k}=\frac{n_{k}}{\sum_{k}^{K}n_{k}}$.
+Client aggregated weight is calculated as $w_{k}=\frac{n_{k}}{\sum_{k}^{K}n_{k}}$.
 $c_{k}$ and $\mathcal{G}$ possess the model $\theta_{k}$ and $\theta_{\mathcal{G}}$ respectively and the models are the same when initialized.
-Learning Objective.
-During the FL training process, each cluster model, on the one hand, is regarded as a mesosphere agent that can optimize the training loss from clients within the cluster; and on the other hand, it serves as a local model to the global model, providing stronger generalization ability.
-In each round of federated training, on the client, it is to learn local model $\theta_{k}$ by minimizing the local empirical loss:
-$F_{i}(\theta_{k})=\sum_{\xi\in\mathcal{D}_{k}}\frac{\ell(\theta_{k},\xi)}{n_{k}}$ (1)
+In each round of federated training, on the client, it is to learn local model $\theta_{k}$ by minimizing the local empirical loss: $F_{i}(\theta_{k})=\sum_{\xi\in\mathcal{D}_{k}}\frac{\ell(\theta_{k},\xi)}{n_{k}}$.
 Where $\ell(\cdot,\cdot)$ is the pre-defined task-driven loss function and $\xi$ represents sampled training data in a batch.
@@
 Therefore, the global objective function can be expressed as:
-$\quad min\quad\sum_{m=1}^{M}\sum_{k\in\mathcal{M}_{m}}w_{k}\ell(\boldsymbol{\theta}_{k})$ (2)
-Model Aggregation. A standard way to obtain a cluster model from clients’ models in CFL is to adopt the FedAvg (Li et al., 2020):
-$\boldsymbol{\Theta}_{m}=\sum_{k\in\mathcal{M}_{m}}w_{k}\boldsymbol{\theta}_{k}$ (3)
-There are various aggregation methods applied to aggregate the global model from clusters’ models.
-FedAdagrad (Reddi et al., 2021b) exploits dynamic aggregation to combine the clients’ parameter updates from the current round with the stored historical parameter updates:
-$\Delta_{m,t}=\boldsymbol{\Theta}_{m,t}-\boldsymbol{\Theta}_{\mathcal{G},t-1}$ (4)
-$\boldsymbol{\Delta}_{t}=\frac{1}{M}\sum_{m=1}^{M}w_{m}\boldsymbol{\Delta}_{m,t}$
-$\boldsymbol{mtumn}_{t}=\beta\boldsymbol{mtumn}_{t-1}+(1-\beta)\Delta_{t},\beta\in[0,1)$
-$\boldsymbol{\upsilon}_{t}=\boldsymbol{\upsilon}_{t-1}+\boldsymbol{\Delta}_{t}^{2}$
-$\boldsymbol{\theta}_{\mathcal{G},t}=\boldsymbol{\theta}_{\mathcal{G},t-1}+\eta\frac{\boldsymbol{mtumn}_{t}}{\sqrt{\boldsymbol{\upsilon}_{t}}+\tau}$
-where $\Delta_{m,t}$ [^1] Both $\Theta_{m,t}$ and $\Theta_{m}$ represent the cluster model.
-Here we add the time stamp $t$ to denote $t$-th round of updating and $\boldsymbol{\Delta}_{t}$ is the difference between the current cluster model and the global model.
-$\boldsymbol{mtumn}_{t}$ is updated iteratively and collects the historical parameter updates, while $\beta$ balances the historical parameter updates $\boldsymbol{mtumn}_{t-1}$ and $\boldsymbol{\Delta}_{t}$.
-A larger $\beta$ indicates the retention of more historical parameter information.
-$\boldsymbol{\upsilon}_{t}$ is an accumulated squared change estimate used to measure changes in parameters and adjust the learning rate to improve the stability and convergence speed of the optimization process.
-MMiC relies on FedAdagrad and derives an extended version (Section 4.3).
-3.2 Banzhaf Power Index
-The Banzhaf Power Index (BPI) is a game theory metric used to quantify the power of an individual voter in a decision-making body where votes are not necessarily equal (Dubey & Shapley, 1979; Laruelle & Valenciano, 2001).
-The index is calculated by considering all possible coalitions (i.e., combinations of voters) that can be formed within a voting body and determining the importance of each voter in each coalition.
-A voter $k$ is considered a core member if their presence in a coalition changes the outcome (denoted as $A_{\mathcal{S}}$) from a loss to a win:
-$k\,is\,core?\begin{cases}True&\text{if }A_{\mathcal{S}^{\prime}}\,is\,a\,loss\text{ and }A_{\mathcal{S}}\,is\,a\,win\\ False&\text{otherwise}\end{cases}$ (5)
-where $\mathcal{S}^{\prime}$ is the set that excludes voter $k$: $\mathcal{S}^{\prime}=\mathcal{S}\setminus\{k\}$.
-BPI could be adopted in federated learning settings to help select the most contributable clients (i.e., a subset of the coalition) during the training process, making the client selection more effective.
-3.3 Markowitz Portfolio Optimzation
-The Markowitz Portfolio Optimization model is an investment optimization model that aims to obtain an optimal investment portfolio with maximized returns and minimized risks (Miller, 1960).
-The model can assess the current investment process through Risk-Adjusted Return Coefficient RARC ($\gamma$).
-The quadratic programming model of Markowitz Portfolio Optimization is formulated as follows:
-$minimize\quad\gamma=\lambda\bar{\sigma}-(1-\lambda)A$ (6)
-indicating RARC is influenced by three variables: investment risk $\bar{\sigma}$, the investment return $A$ and the risk tolerance factor $\lambda$.
-$\lambda$ signifies the investor’s risk tolerance within the range of [0,1] - When $\lambda$ is set to 0.0, the maximum return value is achieved.
-$A$ denotes the total portfolio return across all assets.
-$\bar{\sigma}$ is calculated as follows:
-$\bar{\sigma}=\sum_{i=1}^{K}\sum_{j=1}^{K}w_{i}w_{j}\sigma_{ij}$ (7)
-$\sigma_{ij}=\frac{1}{T-1}\sum_{t=1}^{T}(\alpha_{i,t}-\bar{\alpha}_{i,t-1})(\alpha_{j,t}-\bar{\alpha}_{j,t-1})$
-where $\sigma_{ij}$ is the covariance between assets $i$ and $j$.
-$w_{i}$ and $w_{j}$ are the weights of assets $i$ and $j$, respectively.
-$K$ is the number of assets.
-$\bar{\alpha}_{i,t-1}$ denotes the historical average return of assets $i$ across the preceding $t-1$ rounds.
-$\sigma_{ij}$, the risk between assets $i$ and $j$ is obtained by subtracting the corresponding average return from each round in the history of the asset.
-This formulation can be extended to federated learning scenarios, where $\sigma_{ij}$ quantifies the risk or divergence of local models, and $\bar{\alpha_{i,t-1}}$ reflect the historical performance of individual participants in the federated system.
-4 Methodology
-Figure 1: The MMiC architecture is divided into three layers: Server, Cluster, and Client.
+(1) $\quad min\quad\sum_{m=1}^{M}\sum_{k\in\mathcal{M}_{m}}w_{k}\ell(\bm{\theta}_{k})$
+Model Aggregation.
+We obtain cluster models from clients’ models by using FedAvg (Li et al., 2020): $\bm{\Theta}_{m}=\sum_{k\in\mathcal{M}_{m}}w_{k}\bm{\theta}_{k}$.
+Our method for aggregating cluster models into the global model is elaborated in Section 4.3.
+4. Methodology
+Figure 1.
+The MMiC architecture is divided into three layers: Server, Cluster, and Client.
 The right part illustrates the process of parameter substitution within a cluster.
 The MPO module primarily operates during the aggregation process from clusters to the global model, while the Banzhaf client selection is applied to the client selection within each cluster.
+[…]
 Figure 1 depicts the architecture of MMiC. The left part illustrates the overall architecture including the clients at the bottom, clusters in the middle, and the global server at the top.
 Each cluster accommodates multiple clients, and each client hosts multimodal data with missing modalities locally.
-The cluster model is derived from the federated average aggregation (Section 3.1).
+The cluster model is derived from the federated average aggregation (Section 3).
 Within each cluster, MMiC handles the missing modality issue and the client selection task.
@@
 Second, client selection is performed leveraging the Banzhaf Power Index, which identifies a set of core clients that contribute the most effectively to the cluster’s model (Section 4.2).
-For global aggregation, MMiC extends FedAdagrad by adopting the Markowitz Portfolio Optimization (Section 4.3).
-This method incorporates client training evaluation results from each cluster and reduces the impact of missing modalities by down-weighing the contribution from clusters that contain missing modalities.
-4.1 Cluster-based Parameter Substituion
-We are targeting the multimodal scenario when the dataset has both textual $X^{text}$ and image $X^{image}$ data, thus we form the input as a pair {$X^{text}_{i},X^{image}_{i}$} to the $i$-th data in $\mathcal{D}$.
-The missing modality randomly occurs resulting in data samples in the form of {$X^{null}_{i},X^{image}_{i}$} or {$X^{text}_{i},X^{null}_{i}$} [^2] Please note that MMiC is modality-agnostic – it remains effective regardless of the type or number of modalities, as long as modalities are consistent across clients.
+For global aggregation, MMiC extends $FedAdagrad$ (Reddi et al., 2021a) by adopting the Markowitz Portfolio Optimization (Section 4.3).
+This method incorporates client training evaluation results from each cluster and reduces the impact of missing modalities by down-weighting the contribution from clusters that contain missing modalities.
+4.1. Cluster-based Parameter Substituion
+We consider a multimodal setting where $i$-th data instance is a pair $\{X^{text}_{i},X^{image}_{i}\}$.
+Missing modalities occur randomly, yielding samples such as $\{X^{null}_{i},X^{image}_{i}\}$ or $\{X^{text}_{i},X^{null}_{i}\}$ [^1] Please note that MMiC is modality-agnostic – it remains effective regardless of the type or number of modalities, as long as modalities are consistent across clients.
 We have additional experiments on other modalities shown in Section 5.4.
@@
 Poverty Parameter Identification.
-In one round of client model training, we identify the layer that is the most affected by missing modalities as $poverty$ layer.
-We use the parameter difference between the last batch $b-1$ and the current batch $b$ to define the degree of this effect when the missing modality occurs in batch $b$ and batch $b-1$ is with complete modality [^3] We are aware of the missing modality status of the data.:
-$\Delta\boldsymbol{\theta}_{k,t}^{e,b}(p)$ $=\text{mean}\left(\left|\frac{\boldsymbol{\theta}_{k,t}^{e,b}(p)-\boldsymbol{\theta}_{k,t}^{e,b-1}(p)}{\boldsymbol{\theta}_{k,t}^{e,b-1}(p)}\right|\right)$
-$\text{where}\quad\Delta\boldsymbol{\theta}_{k,t}^{e,b}(p)\in\mathbb{R}$ (8)
-where $e$ is the epoch, $t$ is the training round, $k$ is the client.
-$\Delta\boldsymbol{\theta}_{k,t}^{e,b}(p)$ represents average parameters change rate of the $p$-th layer.
-As $\Delta\boldsymbol{\theta}_{k,t}^{e,b}$ is a set that contains all the layers change rate.
+In one round of client model training, we identify the layer that is the most affected by missing modalities as povertypoverty layer.
+We use the parameter difference between the last batch bb-1 and the current batch $b$ to define the degree of this effect when the missing modality occurs in batch $b$ and batch b1b-1 is with complete modality [^2] We are aware of the missing modality status of the data.:
+(2) $\rho_{k,t}^{e,b}(p)=mean\left(\left|\frac{\bm{\theta}_{k,t}^{e,b}(p)-\bm{\theta}_{k,t}^{e,b-1}(p)}{\bm{\theta}_{k,t}^{e,b-1}(p)}\right|\right)\;\text{and}\;\rho_{k,t}^{e,b}(p)\in\mathbb{R}$
+where $e$ represents the epoch, $t$ is the training round, $k$ denote the client.
+$\rho_{k,t}^{e,b}(p)$ represents average parameter change rate of the $p$-th layer.
+As $\rho_{k,t}^{e,b}$ is a set containing all layers’ change rates.
 We collect all the batches in the $e$-th epoch that contain complete modality data, while the preceding batch contains missing modality data, and form a batch set $\mathcal{B}*$.
 Therefore, the relative change rate for the layer $p$ in one epoch could be expressed as:
-$\Delta\boldsymbol{\theta}_{k,t}^{e}(p)=\frac{\sum_{b\in\mathcal{B}^{*}}\Delta\boldsymbol{\theta}_{k,t}^{e,b}(p)}{|\mathcal{B}^{*}|}$ (9)
-After calculating the average rate of change for all layers in the model in one epoch, we identify the layer with the highest rate of change across all epochs.
-This layer would be considered as the $poverty$ layer with index $\mathcal{P}_{k,t}$:
-$\mathcal{P}_{k,t}=\arg\max_{\begin{subarray}{c}p\in\{1,2,\dots,P\}\end{subarray}}\Delta\boldsymbol{\theta}_{k,t}(p)$ (10)
+(3) $\rho_{k,t}^{e}(p)=\frac{\sum_{b\in\mathcal{B}^{*}}\rho_{k,t}^{e,b}(p)}{|\mathcal{B}^{*}|}$
+After calculating the average change rate for all layers in the model in one epoch, we identify the layer with the highest change rate across all epochs.
+This layer would be considered as the povertypoverty layer with index $\mathcal{P}_{k,t}$:
+(4) $\mathcal{P}_{k,t}=\arg\max_{\begin{subarray}{c}p\in\{1,2,\dots,P\}\end{subarray}}\rho_{k,t}(p)$
 It is worth mentioning that we select parameters from only one layer to reduce communication consumption.
@@
 After identifying the poverty parameter in $c_{k}$, we propose to replace these parameters with the parameters from other clients in the cluster that have complete modality data.
+The proposed parameter substitution can be considered as a model editing strategy (Li et al., 2024).
 Specifically, we compute a weighted average of the corresponding parameters of other clients and replace the poverty parameter in $c_{k}$.
-$\boldsymbol{\theta}_{k,t}(\mathcal{P}_{k,t})$ $=\sum_{j=1}^{J}w_{j}\boldsymbol{\theta}_{j,t}(\mathcal{P}_{k,t})$
-$\quad\text{where }\mathcal{P}_{j,t}=\varnothing,\quad K+J=|\mathcal{M}_{m}|$ (11)
-$c_{k}$ represents the poverty client and $c_{j}$ is the client that trained with the complete modality round, and both $c_{k}$ and $c_{j}$ belong to the cluster $\mathcal{M}_{m}$.
+(5) $\bm{\theta}_{k,t}(\mathcal{P}_{k,t})=\sum_{j=1}^{J}{w_{j}}\bm{\theta}_{j,t}(\mathcal{P}_{k,t})\quad where\quad\mathcal{P}_{j,t}=\varnothing,K+J=|\mathcal{M}_{m}|$
+where $c_{k}$ represents the poverty client and $c_{j}$ is the client that trained with the complete modality round, and both $c_{k}$ and $c_{j}$ belong to the cluster $\mathcal{M}_{m}$.
 Optimizers Could Approximate Parameter Difference.
@@
 However, we can roughly deduce the rate of change by utilizing the gradient derivation function of the optimizer.
-Below, we provide the calculation process for parameter variations during the SGD and Adam optimization processes.
-For SGD, $\eta$ is the learning rate and $\boldsymbol{g}_{k,b}$ is the gradient for $c_{k}$ in $b$ batch.
-$\Delta\boldsymbol{\theta}_{k,t}^{e,b}=\frac{(\boldsymbol{\theta}_{k,t}^{e,b-1}-\eta_{k,t}\boldsymbol{g}_{k,b})-\boldsymbol{\theta}_{k,t}^{e,b-1}}{\boldsymbol{\theta}_{k,t}^{e,b-1}}\\ =\frac{-\eta_{k,t}\boldsymbol{g}_{k,b}}{\boldsymbol{\theta}_{k,t}^{e,b}+\boldsymbol{g}_{k,b}}$ (12)
-Similarly, we can utilize the Adam optimizer to deduce the average rate of change, here $\hat{\boldsymbol{momentum}}$ is the bias-corrected moment estimates and $\hat{\boldsymbol{\nu}}$ is the second-moment estimate in Adam optimizer (Kingma & Ba, 2015):
-$\Delta\boldsymbol{\theta}_{k,t}^{e,b}=\frac{\eta_{k,t}abs(\-\hat{\boldsymbol{momentum}}_{b})}{abs(\boldsymbol{\theta}_{k,t}^{e,b})(\sqrt{\hat{\boldsymbol{\nu}}_{b}}+\epsilon)}$ (13)
-4.2 Banzhaf Client Selection in Clusters
+Below, we provide the calculation process for parameter variations during the Adam optimization processes.
+Here $\hat{\bm{momentum}}$ is the bias-corrected moment estimate and $\hat{\bm{\nu}}$ is the second-moment estimate in Adam optimizer (Kingma and Ba, 2015):
+(6) $\Delta\bm{\theta}_{k,t}^{e,b}=\frac{\eta_{k,t}abs(\-\hat{\bm{momentum}}_{b})}{abs(\bm{\theta}_{k,t}^{e,b})(\sqrt{\hat{\bm{\nu}}_{b}}+\epsilon)}$
+4.2. Banzhaf Client Selection in Clusters
 The standard procedure for current CFL involves randomly selecting a subset of clients $\mathcal{S}_{m}$ from each cluster for training in each round, followed by a global aggregation.
@@
 Few CFL work addresses this issue and quantifies the impact of client selection for each cluster.
-We propose a novel method to fill the gap by leveraging Banzhaf Power Index (BPI) to select the core clients.
+We propose a novel method to fill the gap by leveraging Banzhaf Power Index (Dubey and Shapley, 1979; Laruelle and Valenciano, 2001) (BPI) to select the core clients.
 It comprises two main steps.
@@
 At training round $t$, each cluster $\mathcal{M}_{m}$ makes a random client selection.
-We can obtain the $i$-th client’s local model performance $a_{i,t}$ (i.e., accuracy, F1 score or RSum) of each client in the $t$-th round, and the performance $a_{i,t-1}$ in the ($t$-1)-th round.
+We can obtain the $i$-th client’s local model performance $a_{i,t}$ (i.e., accuracy, F1 score or RSum) of each client in the $t$-th round, and the performance $a_{i,t-1}$ in the (tt-1)-th round.
 We calculate the return $\alpha_{i,t}$ of the client for the $t$-th round based on the difference between these two rounds of performances as follows:
-$\alpha_{i,t}=a_{i,t}-a_{i,t-1}$ (14)
+(7) $\alpha_{i,t}=a_{i,t}-a_{i,t-1}$
 where $\alpha_{i,t}<0$ indicates that the client is generating a negative return to the investment portfolio; conversely, $\alpha_{i,t}>0$ signifies a positive return.
 Then, the cluster’s return is calculated as the sum of the client returns:
-$A_{\mathcal{S}_{m,t}}=\sum_{i=0}^{|\mathcal{S}_{m,t}|}w_{i}\alpha_{i,t}$ (15)
-For each cluster $\mathcal{S}_{m,t}$, we have its total return $A_{\mathcal{S}_{m,t}}$ by Equation (15).
-Correspondingly, we can calculate the total return for $\mathcal{S}_{m,t}^{{}^{\prime}}$ when excluding client $c_{i}$, $\𝑖\mathcal{S}_{m,t}^{\prime}=\mathcal{S}_{m,t}\textbackslash\{i\}$: Then, we use the return of cluster model to evaluate the core members:
-$c_{i}\,is\,core\,member?\begin{cases}True&A_{\mathcal{S}_{m,t}}\geq\alpha^{m}_{t}\text{ and }A_{\mathcal{S}_{m,t}^{{}^{\prime}}}<\alpha^{m}_{t}\\ False&\text{otherwise}\end{cases}$ (16)
+(8) $A_{\mathcal{S}_{m,t}}=\sum_{i=0}^{|\mathcal{S}_{m,t}|}w_{i}\alpha_{i,t}$
+[…]
+(9) $c_{i}\,is\,core\,member?\begin{cases}True&A_{\mathcal{S}_{m,t}}\geq\alpha^{m}_{t}\text{ and }A_{\mathcal{S}_{m,t}^{{}^{\prime}}}<\alpha^{m}_{t}\\ False&\text{otherwise}\end{cases}$
 The above equation describes the process of determining whether client $c_{i}$ is a core member during the $t$-th round of client random selection.
 We utilize a counter $\phi(*)$ that starts with 0 to tally the number of times each client has been identified as a core member:
-$\phi(i)=\sum_{t=1}^{T}\mathbb{I}\left(c_{i}\,is\,a\,core\,member\,\right)$ (17)
+(10) $\phi(i)=\sum_{t=1}^{T}\mathbb{I}\left(c_{i}\,is\,a\,core\,member\,\right)$
 where $\mathbb{I}(*)$ is the indicator function that returns 1 if the condition * is satisfied, otherwise returns 0.
 Client Selection Probability.
+BPI prioritizes clients without missing modalities, thus favoring those that contribute more to performance improvements.
 We could design a method to measure the probability of a client being selected in the current round based on its core times.
@@
 To solve this problem, we set another counter $\mathcal{T}(*)$ that records the number of times the client has been selected in its history, and we determine the final selection probability by balancing the historically selected times $\mathcal{T}(i)$ and the core times $\phi(i)$:
-$prob(i)=\frac{e^{\frac{\tau\phi(i)}{\mathcal{T}(i)}}}{\sum_{i\in\mathcal{M}_{m}}e^{\frac{\tau\phi(i)}{\mathcal{T}(i)}}}$ (18)
+(11) $prob(i)$ $=\frac{e^{\frac{\tau\phi(i)}{\mathcal{T}(i)}}}{\sum_{i\in\mathcal{M}_{m}}e^{\frac{\tau\phi(i)}{\mathcal{T}(i)}}}$
 For client $i$ in cluster $\mathcal{M}_{m}$, its probability in the $t$-th round to be selected is $prob(i)$.
@@
 As $\tau$ increases, the disparity in probability increases, and as $\tau$ decreases, the disparity in probability decreases.
-4.3 Markowitz Profolio Optimization for Global Aggregation
-As introduced in Section 3.3, Markowitz Portfolio Optimization is a framework for constructing an investment portfolio that aims to maximize returns for a given level of risk.
-We are the first to integrate this theory into CFL by considering to mitigate the negative effects of local client models (e.g., those with missing modalities) on the global model.
-Specifically, we propose a mechanism where, in each communication round, the performance of the selected clients within a cluster $\mathcal{M}_{m}$ is quantified through a non-linear scoring function.
-This score serves as a direct indicator of the cluster’s performance during the round, allowing for adaptive control over the magnitude of the global model’s updates.
-By leveraging this approach, we dynamically regulate the contribution of each cluster to the global model aggregation, thereby reducing the impact of underperforming or unreliable clients while maintaining the global model’s robustness.
-To achieve this, we require three core indicators to construct our “investment portfolio” for the cluster $\mathcal{M}m$: the return (i.e., payoff) of each cluster ($A_{m,t}$, Equation (15)); the risk associated with the cluster ($\bar{\sigma}_{m,t}$); and the tolerance for return-risk ratio for each cluster ($\lambda$).
-Cluster Risk. According to the asset risk calculation formula Equation (7), we obtain the covariance as follows:
-$\bar{\sigma}_{m,t}$ $=\sum_{i=1}^{|\mathcal{S}_{m}|}\sum_{j=1}^{|\mathcal{S}_{m}|}w_{i}w_{j}\Bigg{\{}\frac{1}{T-1}\sum_{t=1}^{T}$
-$\quad(\alpha_{i,t}-\bar{\alpha}_{i,t-1})(\alpha_{j,t}-\bar{\alpha}_{j,t-1})\Bigg{\}}$ (19)
-where $\bar{\alpha}_{i,t-1}$ represents the historical average return for client $i$ in round $t-1$.
-$w_{i}$ and $w_{j}$ are weights assigned to the $i$-th and $j$-th clients respectively.
-Please refer to Section 3.1 for weights’ calculation.
-Then the cluster risk is formulated as follows according to Equation (7):
-$\quad\gamma_{m,t}=\lambda\bar{\sigma}_{m,t}-(1-\lambda)A_{m,t}$ (20)
+4.3. Markowitz Portfolio Optimization for Global Aggregation
+In our framework, MMiC builds upon $FedAdagrad$ (Reddi et al., 2021a) to support adaptive cluster-level aggregation in the presence of modality incompleteness.
+To enhance global robustness, we propose a novel integration of Markowitz Portfolio Optimization (MPO) into $FedAdagrad$.
+Originally introduced in financial theory (Miller, 1960), MPO constructs investment portfolios that trade off expected return and associated risk.
+Inspired by this principle, we adapt the concept to dynamically regulate the influence of each cluster in global aggregation.
+Specifically, clusters consisting of underperforming clients—such as those suffering from missing modalities—may introduce biased or unreliable updates.
+We treat each cluster as an “investment” and aim to down-weight risky clusters while preserving beneficial contributions.
+Our approach introduces a scoring function that quantifies the overall quality of a cluster’s update in each round, based on both its expected return and internal risk.
+This score directly informs the global aggregation process, enabling us to balance between current update quality and accumulated historical knowledge.
+To achieve this, we require three core indicators to construct our “investment portfolio” for cluster $\mathcal{M}m$ includes the expected return (i.e., payoff) of the cluster, denoted as $A_{m,t}$ (Equation 8); the risk $\bar{\sigma}_{m,t}$, derived from the intra-cluster covariance; and the return-risk tolerance coefficient $\lambda$ that governs the trade-off between return and risk for the cluster.
+Cluster Risk. According to the asset risk calculation formula from MPO (Miller, 1960; Markowitz, 1991), we obtain the covariance as follows:
+(12) $\bar{\sigma}_{m,t}=\sum_{i=1}^{|\mathcal{S}_{m}|}\sum_{j=1}^{|\mathcal{S}_{m}|}w_{i}w_{j}\left\{\frac{1}{T-1}\sum_{t=1}^{T}(\alpha_{i,t}-\bar{\alpha}_{i,t-1})(\alpha_{j,t}-\bar{\alpha}_{j,t-1})\right\}$
+where $\bar{\alpha}_{i,t-1}$ represents the historical average return for client $i$ in round t1t-1.
+$w_{i}$ and $w_{j}$ are weights assigned to the $i$-th and $j$-th clients respectively (Please refer to Section 3 for weights’ calculation).
+Then the Risk-Adjusted Return Coefficient (RARC) is formulated as follows:
+(13) $\quad\gamma_{m,t}=\lambda\bar{\sigma}_{m,t}-(1-\lambda)A_{m,t}$
 where $\lambda$ [0,1] is the risk tolerance constant.
 We defaultly set it as 0.5 to balance risk and return.
-Our goal is to minimize the sum of weighted cluster RARC (cluster weight calculation is introduced in Section 3.1):
-$minimize\quad\gamma_{t}=\sum_{m=1}^{M}w_{m}\gamma_{m,t}$ (21)
-We analyzed how we achieved the time complexity of MPO in Appendix D.
+Section 5.8 demonstrates how we reduce the cubic computational complexity of MPO to a quadratic complexity.
 Participate in Global Aggregation.
-A higher $\gamma_{t}$ signifies that the clusters have undergone a relatively poor local client training process.
-In such scenarios, our objective is to diminish the impact of the current round’s parameter updates on the global aggregation by down-weighting the cluster parameters associated with higher $\gamma$.
-To accomplish this, we introduce a dynamic $\beta^{*}$ to replace the static $\beta$ in FedAdagrad, resulting in FedAdagradMPO, which effectively captures the influence of $\gamma_{t}$:
-$\beta^{*}=\beta+\frac{tanh(ReLU(\gamma_{t}))}{\tau}$ (22)
-When $\gamma_{t}>0$, then $\beta^{*}>\beta$, which helps retain more historical parameter updates in the global aggregation.
-Otherwise, $\beta^{*}=\beta$.
+We have evaluated the training RARC $\gamma_{m,t}$ for each cluster.
+Next, we dynamically incorporate these RARC into the global aggregation process.
+A higher $\gamma_{m,t}$ indicates that the cluster has experienced a relatively poor local training process.
+In such scenarios, our objective is to diminish the impact of the current round’s parameter updates on the global aggregation by down-weighting the cluster parameters associated with higher $\gamma_{m,t}$.
+To accomplish dynamic aggregation between current round and historical rounds, $FedAdagrad$ employs the following update mechanism: $\bm{\Delta_{m,t}}$ is the difference between the current cluster model and the global model.
+$\bm{mtn}_{m,t}$ is updated iteratively and collects the historical parameter updates, while $\beta$ balances the historical parameter updates $\bm{mtn}_{m,t-1}$ and $\bm{\Delta}_{m,t}$.
+A larger $\beta$ indicates the retention of more historical parameter information.
+$\bm{\upsilon}_{t}$ is an accumulated squared change estimate used to measure changes in parameters and adjust the learning rate to improve the stability and convergence speed of the optimization process:
+(14) $\bm{\Delta_{m,t}}=w_{m}(\bm{\Theta}_{m,t}-\bm{\Theta}_{\mathcal{G},t-1})$
+$\bm{mtn}_{m,t}=\beta\bm{mtn}_{m,t-1}+(1-\beta)\bm{\Delta_{m,t}},\beta\in[0,1)$
+$\bm{\upsilon}_{m,t}=\bm{\upsilon}_{m,t-1}+\bm{\Delta}_{m,t}^{2}$
+$\bm{\theta}_{\mathcal{G},t}=\bm{\theta}_{\mathcal{G},t-1}+\eta\frac{\bm{mtn}_{m,t}}{\sqrt{\bm{\upsilon}_{m,t}}+\tau}$
+We introduce a dynamic $\beta_{m,t}^{*}$ to replace the static $\beta$ in $FedAdagrad$, which effectively captures the influence of $\gamma_{m,t}$:
+(15) $\beta_{m,t}^{*}=\beta+\frac{tanh(ReLU(\gamma_{m,t}))}{\tau}$
+$\bm{mtn}_{m,t}=\beta_{m,t}^{*}\bm{mtn}_{m,t-1}+(1-\beta_{m,t}^{*})\bm{\Delta_{m,t}},\beta_{m,t}^{*}\in[0,1)$
+When $\gamma_{m,t}>0$, then $\beta_{m,t}^{*}>\beta$, which helps retain more historical parameter updates in the global aggregation.
+Otherwise, $\beta_{m,t}^{*}=\beta$.
 Here $\tau$ serves as the smoothness control constant for the range.
 We suggest setting the temperature to range [10,20], and it depends on the value set for $\beta$.
-4.4 Variants of MMiC
+4.4. Variants of MMiC
 We propose two MMiC variants by extending two existing methods that address modality incompleteness.
@@
 In this way, the missing modality information could be filled in the shared space.
-5 Experiments
-Table 2: Global and personalized performance compared with different algorithms.
+5. Experiments
+5.1. Experiment Setup
+Table 1.
+Global and personalized performance compared with different algorithms.
 Bold with * indicated the best performance, while bold black indicated excellent or stable top three performance.
@@
 MMiC (GF) 19.625 ± 4.797 *22.796 ± 6.04 *18.450 ± 2.73 *84.715 ± 3.13 83.639 ± 2.10 *84.601 ± 4.69 *112.506 ± 5.10 *160.99 ± 6.082 *124.414 ± 2.86
-5.1 Datasets, Models and Metrics
+5.1.1. Datasets, Models and Metrics
 Our evaluation of MMiC is based on two multimodal classification datasets and one multimodal retrieval dataset.
-Because each dataset has distinct characteristics, we followed (Feng et al., 2023; Jeni et al., 2013) in employing different base models and evaluation metrics for each dataset.
-We also provide learning configurations in Appendix A.
+Because each dataset has distinct characteristics, we followed (Feng et al., 2023; Jeni et al., 2013) in employing different base models, learning configurations and evaluation metrics for each dataset.
 CrisisMMD (Ofli et al., 2020).
@@
 We brought in a lightweight multi-modal model that used pre-trained embedding layers for input and concatenates shared representations (Gupta et al., 2024; Feng et al., 2023).
-Following (Feng et al., 2023), we employed the F1 (Goutte & Gaussier, 2005) score to evaluate the performance of CrisisMMD.
+Following (Feng et al., 2023), we employed the F1 (Goutte and Gaussier, 2005) score to evaluate the performance of CrisisMMD.
 UPMC-Food101 (Gallo et al., 2020).
-This dataset was sourced from Food101, containing 101,000 image-text pairs across 101 food categories.
+UPMC-Food101 sourced from Food101, containing 101,000 image-text pairs across 101 food categories.
 We used CLIP with the official ViT version of the backbone pre-trained model (Radford et al., 2021).
@@
 We relied on the RSum as the evaluation metric to measure the performance of the retrieval task (Faghri et al., 2018).
-5.2 Benchmark Algorithms
+Learning settings.
+In all of our experiments, we set a global learning rate (0.1, 0.01, 0.05) in the three datasets (CrisisMMD, Food101, Flickr30k) respectively to global optimizers (both $FedAdagrad$ and $FedAdagrad_{MPO}$).
+Our Parameter difference calculation process is derived from Equation 6 and we used the Adam optimizer with a learning rate of 0.005 for the clients.
+These parameters were obtained through our simple tuning in the early stage of our experiment.
+The batch size for the CrisisMMD dataset was set to 15, and 7 for the UPMC-Food101 dataset, 20 for the Flickr30k dataset.
+All the local epoch was set to 3.
+The loss function on the client was cross-entropy, except for MMiC (GMC-Fusion), where the loss consists of two parts: cross-entropy and contrastive loss between existing modalities (Poklukar et al., 2022).
+5.1.2. Benchmark Algorithms
 We compared MMiC with selected representative works and state-of-the-art works.
-FedAvg (McMahan et al., 2017) represented a basic federated architecture where multiple clients collaborated with a central server.
-FedOpt (Reddi et al., 2021a) proposed three specialized federated aggregation methods, building upon FedAvg.
-FedSoft (Ruan & Joe-Wong, 2022), as a soft clustering method, introduced a technique that permitted a client to belong to multiple clusters.
-PACFL (Vahidian et al., 2023), an advanced CFL methodology, employed Singular Value Decomposition (SVD) to deconstruct the dataset into its constituent subspaces, subsequently applying hierarchical clustering techniques to these subspaces for effective clustering.
-CreamFL (Yu et al., 2023), was a MFL algorithm that employed inter-modal contrastive learning to address the issue of modality imbalance.
-PmcmFL (Bao et al., 2024), was another MFL for addressing missing modality.
-They utilized prototype aggregation to address the issue of missing modalities in classification tasks.
-5.3 Settings
-Cluster and Clients Settings.
+FedAvg (McMahan et al., 2017) (first work for FL), FedOpt (Reddi et al., 2021b) (outstanding FL work), FedSoft (Ruan and Joe-Wong, 2022) (a soft SOTA clustered FL), PACFL (Vahidian et al., 2023) (an advanced CFL methodology), CreamFL (Yu et al., 2023) (MFL for addressing missing modality), and PmcmFL (Bao et al., 2024) (a SOTA work for addressing missing modality).
+5.1.3. Cluster and Clients Settings
 In this experiment, we employed two distance metrics for clustering: Locality Sensitive Hashing (LSH) sketch (Yang et al., 2024; Liu et al., 2023) and SVD decomposition (Vahidian et al., 2023).
@@
 Additionally, the global server provided accessibility to the test set for all clients.
-We simulated the Non-IID scenario by assigning a label probability table to each client utilizing the Dirichlet distribution (Teh, 2010) with $\beta_{diri}=0.5$.
-For all the experiments, we set the number of clients to 50 so that ensured each client had at least 200 training data.
+We simulated the Non-IID scenario by utilizing the Dirichlet distribution (Teh, 2010) with $\beta_{diri}=0.5$.
+For all the experiments, we set the number of clients to 50 so that ensured each client had at least 200 training samples.
 Our client selection strategy comprised an initial phase of 100 rounds with random selection, followed by a subsequent phase of 200 rounds utilizing Banzhaf client selection (300 global rounds).
-Missing Modality Settings.
+5.1.4. Missing Modality Settings.
 Our setting with missing modalities was governed by two variables: clients with missing modality (mc) and missing modality rate (mm), denoting the proportion of all clients selected to have missing modalities and the proportion of missing samples on these clients, respectively.
 Our main experiment (Section 5.2) offered three levels of modality incompleteness settings: (mm@0.2, mc@0.3), (mm@0.2, mc@0.5) and (mm@0.5, mc@0.5).
+(mm@0.2, mc@0.3) means ratio of mm = 20% and ratio of mc = 30%.
 Each client has a fixed number of missing samples, determined by the mm rate.
@@
 Since the selected clients use a random training subset for training in each round, the maximum missing modality rate for a client in that round is $min(\frac{missing\ samples}{subset\ samples},1)$.
-5.4 Overall Performance
+5.2. Overall Performance
 We evaluated MMiC based on global performance (gbp) and personalized performance (plp) and presented the results in Table 1.
@@
 PmcmFL currently supports only classification tasks; however, its performance in classification tasks is somewhat limited.
-(a) only PPS
-(b) w/o PPS
-(c) only BPI
-(d) w/o BPI
-(e) only MPO
-(f) w/o MPO
-Figure 2: Ablation study under CrisisMMD (mm@0.5 mc@0.5).
-5.5 Ablation Study
-We conducted ablation experiments of MMiC on CrisisMMD.
-The results are shown in Figure 2.
-A further ablation study on Food101 and Flickr30k is provided in Appendix B.
-BPI (b), MPO (p) and PPS (s) represent Banzhaf Power Selection, Markowitz Portfolio Optimization, and Poverty Parameter Substitution, respectively.
-The figures show that MMiC consistently achieves higher accuracy in the final rounds across multiple ablation tests.
-Specifically, both ‘w/o PPS’ (Figure 2(b)) and ‘w/o BPI’ (Figure 2(d)) saw a decrease of more than 5% compared with MMiC.
-The final performance of ‘w/o MPO’ (Figure 2(f)) is close to MMiC, yet it remained approximately 2% lower than MMiC.
-In addition, removing any of the two contributions leads to varying degrees of performance decline, with the resulting performance consistently falling between that of MMiC and the configuration with a single module removed (Figure 2(a), 2(c), 2(e)).
+Overall, MMiC consistently outperforms traditional methods for addressing missing modality challenges in FL.
+5.3. Ablation Study
+We conducted an ablation study on the CrisisMMD, Food101, and Flickr30k datasets under the settings of mm@0.2 and mc@0.5 (Table 2).
+The ablation modules are denoted as ‘b’, ‘p’, and ‘s’ (BPI, MPO and PPS respectively).
+It shows that MMiC consistently achieves higher accuracy in the final rounds across multiple ablation tests.
+Specifically, ‘w/o b’ and ‘w/o s’ saw a decrease of more than 5% compared to MMiC in CrisisMMD while others still performed poorly.
+The final performance of ‘w/o p’ is close to MMiC, yet it remained approximately 2%, 0.2% and 3% lower than MMiC in three datasets respectively.
+In addition, removing any of the two contributions leads to varying degrees of performance decline, with the resulting performance consistently falling between that of MMiC and the configuration with a single module removed in Food101 and Flickr30k.
 We conclude that removing any module leads to performance degradation.
-We further proved the effectiveness of BPI in Appendix C.
-5.6 Experiments on other modalities
-We further experimented on two additional multimodal datasets with modalities other than image and text, aiming to show that MMiC is modality-agnostic.
-UCI-HAR (Reyes-Ortiz et al., 2013) and KU-HAR (Sikder & Nahid, 2021) both encompass smartphone sensor data utilized for the classification of diverse human activities.
+We further proved the effectiveness of BPI in Section 5.5 and MPO in Section 5.6.
+We believe that BPI helps MMiC by reducing the weight of clients heavily affected by missing modalities, ensuring well-performing clients are prioritized.
+MPO further improves aggregation by minimizing the influence of high-risk clusters with missing modalities.
+Table 2. Ablation study on CrisisMMD, Food101 and Flickr30k under (mm@0.2, mc@0.5).
+Configuration CrisisMMD (Acc) Food101 (Acc) Flickr30k (RSum)
+w/o (b, p, s) 57.22 86.97 4.28
+w/o s 58.91 88.54 6.99
+w/o p 61.07 88.91 8.18
+w/o b 57.03 87.26 8.10
+only s 60.36 88.03 7.08
+only p 59.87 87.49 4.84
+only b 61.24 87.41 5.36
+MMiC 63.55 90.12 11.02
+5.4. Experiments on other modalities
+In this section, we further broaden our experimental evaluation by incorporating more modalities, aiming to investigate the feasibility of our algorithm on non-dominant modalities, aiming to show that MMiC is modality-agnostic.
+UCI-HAR (Rasekh et al., 2014) and KU-HAR (Sikder and Nahid, 2021) both encompass smartphone sensor data utilized for the classification of diverse human activities.
 They include accelerometer and gyroscope readings gathered from a multitude of participants.
 We compared MMiC with FedOpt, which exhibited the good performance in Section 5.2 , for this analysis.
-The experimental setup involved 30 clients running 150 global rounds, with all other parameters remaining consistent with those outlined in Section 5.3.
+The experimental setup involved 30 clients running 150 global rounds, with all other parameters remaining consistent with those outlined in Section 5.1.3.
 The findings of this experiment are presented in Table 3.
 MMiC demonstrated a pronounced superiority over FedOpt, with enhancements of 2% and 6% in the KU-HAR and UCI-HAR respectively.
-Table 3: Performance comparison of MMiC and FedOpt on UCI-HAR and KU-HAR datasets.
+Table 3. Performance comparison of MMiC and FedOpt on UCI-HAR and KU-HAR datasets.
 Dataset/Algorithm MMiC (%) FedOpt (%)
-UCI-HAR 86.99 80.89
-KU-HAR 75.83 73.24
-6 Conclusions
-In this work, we introduce MMiC, a model-independent and modality-agnostic framework designed to mitigate the impact of missing modalities in multimodal federated learning.
-By substituting parameters within clustered federated learning, adapting the Banzhaf Power Index for client selection, and extending Markowitz Portfolio Optimization for global aggregation, MMiC demonstrates improved performance over existing federated architectures in addressing missing modalities.
-Moving forward, we plan to further enhance the framework’s robustness by substituting additional poverty parameters and exploring more efficient solutions.
-Appendix A Learning Settings
-In all of our experiments, we set a global learning rate (0.1, 0.01, 0.05) in the three datasets (CrisisMMD, Food101, Flickr30k) respectively to global optimizers ($FedAdagrad_{MPO}$, $FedAdagrad$).
-Our Parameter difference calculation process is derived from Equation 13 and we used the Adam optimizer with a learning rate of 0.005 for the clients.
-These parameters were obtained through our simple tuning in the early stage of our experiment.
-The batch size for the CrisisMMD dataset was set to 15, and 7 for the UPMC-Food101 dataset, 20 for the Flickr30k dataset.
-All the local epoch was set to 3.
-The loss function on the client was cross-entropy, except for MMiC (GMC-Fusion), where the loss consists of two parts: cross-entropy and contrastive loss between existing modalities (Poklukar et al., 2022).
-Appendix B Ablation Study on Food101 and Flickr30k
-We conducted an ablation study on the Food101 and Flickr30k datasets under the settings of mm@0.2 and mc@0.5.
-The modules are denoted as ‘b’, ‘p’, and ‘s’ (BPI, MPO and PPS respectively following Section 5.5).
-The full MMiC framework demonstrated robust performance, achieving an accuracy exceeding 90% on Food101 and a score of 11 on Flickr30k under the specified conditions.
-In contrast, the exclusion of all modules lead to a significant decline in accuracy, reducing it to 86% on Food101 and 4.3 on Flickr30k.
-The removal of a single module generally yields superior performance compared to the exclusion of two modules, as presented in Table 4.
-Table 4: Ablation study on Food101 and Flickr30k under (mm@0.2, mc@0.5).
-Configuration Food101 (Accuracy) Flickr30k (RSum)
-w/o (b, p, s) 86.97 4.28
-w/o s 88.54 6.99
-w/o p 88.91 8.18
-w/o b 87.26 8.10
-only s 88.03 7.08
-only p 87.49 4.84
-only b 87.41 5.36
-MMiC 90.12 11.02
-Figure 3: The number of additional training rounds needed for the Random Client Selection to achieve the same performance as the Banzhaf Client Selection (Flickr30k).
-Appendix C Effectiveness of Banzhaf Client Selection
-The federated architecture had shown the feasibility of client selection, as demonstrated in (Nishio & Yonetani, 2019; Cho et al., 2020).
+UCI-HAR (Acc) 86.99 75.83
+KU-HAR (Acc) 80.89 73.24
+Figure 2.
+The number of additional training rounds needed for the Random Client Selection to achieve the same performance as the Banzhaf Client Selection (Flickr30k).
+The blue bar means random selection, and the yellow represents the BPI selection.
+5.5. Effectiveness of Banzhaf Client Selection
+The federated architecture had shown the feasibility of client selection, as demonstrated in (Nishio and Yonetani, 2019; Cho et al., 2020).
+In this section, we compared training efficiency under missing modalities and show that BPI achieves higher efficiency than random selection.
 The Banzhaf client selection strategy aimed to optimize global aggregation efficiency by minimizing communication rounds.
@@
 It should be noted that randomly selected patterns required additional (12, 24, or 16) training rounds to surpass or approach the performance of the Banzhaf client selection.
-Appendix D Time Complexity of MPO
+To some extent, MMiC facilitates faster convergence of the global model under missing modality settings by prioritizing the selection of core clients.
+5.6. MPO: A Potent Solution against Over-fitting
+In this section, we show additional features of MPO against overfitting.
+We only implemented MMiC with MPO and deactivated PPS and BPI.
+Therefore, the comparison is between $FedAdagrad_{MPO}$ (MMiC) and $FedAdagrad$ (FedOpt) on dataset CrisisMMD (mm@0.2, mc@0.3) .
+Our purpose is to evaluate how the dynamic updating of parameters in $FedAdagrad_{MPO}$ contributes to enhanced stability.
+During the training, it had been converged after 400 rounds.
+We continued the training after 400 rounds while observing the personalized performance of $FedAdagrad$ and $FedAdagrad_{MPO}$, which showed the stability of our algorithm.
+The personalized evaluation is based on the average weighted performance of each client.
+It can be observed that the personalized performance of FedOpt begins to decline after round 413 and never recovers to its peak level thereafter.
+This indicated that some clients have already experienced overfitting, leading to a degradation in performance.
+In contrast, although the performance curve of MMiC experiences a short decline after round 543 as shown in Figure 3.
+$FedAdagrad_{MPO}$ forms risk by calculating the covariance of historical multi-round returns between pairs of clients (Equation 12).
+When partial clients overfit, the covariance increases, leading to a higher risk and an overall increase in RARC.
+When the majority of clients overfit, the total return decreases, yet the overall RARC still increases.
+An increase in RARC indicates that fewer parameters will be aggregated to the global model.
+The experimental results demonstrated that the stability of our algorithm is significantly enhanced by incorporating the MPO, particularly in mitigating performance degradation caused by overfitting during the training of clients.
+Figure 3.
+Personalized performance of comparison after continuous training since the model convergence (400 rounds).
+The two dotted lines are the average scores.
+The upper part (blue) represents $FedAdagrad_{MPO}$ (MMiC) and the lower part (yellow) represents FedAdagradFedAdagrad (FedOpt).
+Figure 4.
+The two plots on the left depict the SCV of LSH clustering under two different hyperparameter (number of clusters), while the two plots on the right show the SCV of SVD-based clustering with the threshold hyperparameter.
+The numbers within the images represent the indices of the cluster labels.
+5.7. Global Performance Influenced by Clustered Federated Algorithms
+We explored the impact of clustering quality on MMiC performance.
+Figure 4 plots Silhouette Coefficient Value (SCV) of two clustering algoritms with two different hyperparameters respectively.
+This experiment was conducted on dataset CrisisMMD (mm@0.2, mc@0.5) and we investigated how to change clustering hyperparameters to get a higher SCV.
+The SCV quantifies the ratio of the intra-cluster distance, defined as the distance from the sample point to the nearest cluster to which it doesn’t belong (Shahapure and Nicholas, 2020).
+A larger SCV signifies a greater separation between the cluster structure containing the sample point and its nearest neighbouring cluster structure, thereby indicating superior clustering quality under these conditions.
+For Locality Sensitive Hashing (LSH) clustering, their clustering algorithm is based on KMeans (Kanungo et al., 2002), which requires specifying the number of clusters.
+We adjusted the number of clusters from 1 to 20 to identify the two highest SCV.
+When the number of clusters is set to 3, the SCV attained its peak at 0.185.
+With 7 clusters and its second highest value of 0.125.
+For Singular Value Decomposition (SVD) clustering, which employs hierarchical clustering (Johnson, 1967), we varied the threshold from 1 to 20 to determine the optimal thresholds, identifying two values: 10 (SCV = 0.18) and 14 (SCV = 0.2).
+This preliminary selection of clustering parameters assisted us in identifying better and more suitable clustering algorithms and parameters for MMiC.
+Figure 5 investigated the selection process of our clustering strategies.
+What we found are: (1) It’s shown that Poverty Parameter Substitution within clusters heavily relies on the quality of client clustering.
+For instance, in LSH clustering, a higher clustering quality (a higher SCV) often correlates with better global performance (LSH: 0.185 compared with LSH: 0.126).
+(2) By comparing different clustering methods, we can infer the contribution of each clustering algorithm to MMiC based on global performance.
+It is evident that LSH clustering outperformed SVD 0.2 in terms of global model performance when the SCV reached at 0.185.
+(3) We included manual clustering as a gold standard (clustering based on the overlap of client labels).
+Overall, both LSH (0.185) and SVD (0.2) clustering achieved global model performance comparable to that of manual clustering, making them reliable references for the clustering algorithms employed in this study.
+When compared to SVD, LSH demonstrated slightly superior performance.
+Figure 5. We compared two clustering algorithms, LSH and SVD, by selecting the top-2 SCV hyperparameters for each and evaluating their global performance on MMiC. The blue curve is manual clustering.
+5.8. Time Complexity of MPO
 Since the Banzhaf Client Selection do not incur an additional cost, the time complexity of MMiC compared to other works depended on the complexity of Markowitz Portfolio Optimization.
@@
     $O(\frac{K^{2}}{M})$ was less than $O(\frac{K^{2}\cdot T}{M})$ apparently. Therefore, the time complexity of MPO is $O(\frac{K^{2}}{M}\cdot T+K\cdot T)$.
+6. Conclusions
+In this work, we introduce MMiC, a model-independent and modality-agnostic framework designed to mitigate the impact of missing modalities in multimodal federated learning.
+By substituting parameters within clustered federated learning, adapting the Banzhaf Power Index for client selection, and extending Markowitz Portfolio Optimization for global aggregation, MMiC demonstrates improved performance over existing federated architectures in addressing missing modalities.
+Moving forward, we plan to further enhance the framework’s robustness by substituting additional poverty parameters and exploring more efficient solutions.
+7. GenAI Usage Disclosure
+The authors acknowledge the use of an AI language model to assist in the linguistic refinement of selected sections of this paper, specifically the Introduction and Methodology.
+All the wordings, equations, and conclusions were revised again after being reviewed by the authors.